### v1

# An Approximation Algorithm for Monotone Submodular Cost Allocation

Abstract
In this paper, we consider the minimum submodular cost allocation (MSCA) problem. The input of MSCA is $k$ non-negative submodular functions $f_{1},\ldots,f_{k}$ on the ground set $N$ given by evaluation oracles, and the goal is to partition $N$ into $k$ (possibly empty) sets $X_{1},\ldots,X_{k}$ so that $\sum_{i=1}^{k}f_{i}(X_{i})$ is minimized. In this paper, we focus on the case when $f_{1},\ldots,f_{k}$ are monotone (denoted by Mono-MSCA). We provide a natural LP-relaxation for Mono-MSCA, which is equivalent to the convex program relaxation introduced by Chekuri and Ene. We show that the integrality gap of the LP-relaxation is at most $k/2$, which yields a $k/2$-approximation algorithm for Mono-MSCA. We also show that the integrality gap of the LP-relaxation is at least $k/2-\epsilon$ for any constant $\epsilon>0$ when $k$ is fixed.
1 Introduction
Let $N$ be a finite set of cardinality $n$, and $k\geq 2$ a positive integer. A set function $f:2^{N}\to\mathbb{R}$ is called submodular if $f(S)+f(T)\geq f(S\cup T)+f(S\cap T)$ holds for every $S,T\subseteq N$. In this paper, we focus on the Minimum Submodular Cost Allocation (MSCA) problem, introduced by Chekuri and Ene [5]. The input of MSCA is $k$ non-negative submodular functions $f_{1},\ldots,f_{k}:2^{N}\to\mathbb{R}_{+}$ given by evaluation oracles, and the goal is to partition $N$ into $k$ (possibly empty) sets $X_{1},\ldots,X_{k}$ so that $\sum_{i=1}^{k}f_{i}(X_{i})$ is minimized, i.e.,
$\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$
Throughout this paper, we assume that the input functions $f_{1},\ldots,f_{k}$ satisfy $f_{i}(\emptyset)=0$ for each $i=1,\ldots,k$. Several well-studied problems such as the multiway cut problems in graphs and hypergraphs, the submodular multiway partition problems, and uniform metric labeling can be cast as special cases of MSCA. In the case when all submodular functions are monotone, MSCA is equivalent to the submodular facility location problem considered by Svitkina and Tardos [18]. In this paper, we focus on this setting, i.e., the case where each of $f_{1},\ldots,f_{k}$ is monotone submodular. We denote MSCA with monotone submodular functions $f_{1},\ldots,f_{k}$ by Mono-MSCA. The aim of this paper is to investigate the integrality gap of a natural LP-relaxation for Mono-MSCA when $k$ is small.
LP-relaxation.
To provide an LP-relaxation for Mono-MSCA, we introduce some notation. Let $\mathbf{1}\in\mathbb{Z}^{N}_{>0}$ denote the all-ones vector and $k]:=\{1,2,\ldots,k\}$. For any $S\subseteq N$, let $\chi_{S}\in\{0,1\}^{N}$ denote the characteristic vector of $S$. We consider the following linear programming relaxation for MSCA:
$(\textsc{LP-Rel})\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
$\mathrm{subject\ to}$ $\ \ \ y_{i}\in\mathbb{R}^{2^{N}}_{+}\ \mathrm{for\ every\ }i\in[k],$
$\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)\chi_{S}=\mathbf{1}.$
In this LP-relaxation, we interpret $y_{i}(S)=1$ to mean that the set $S$ is assigned to $f_{i}$ in MSCA, whereas $y_{i}(S)=0$ means that $S$ is not assigned to $f_{i}$. We note that if an integer feasible solution $(y_{1},\ldots,y_{k})$ satisfies $y_{i}(S_{1})=y_{i}(S_{2})=1$ for two distinct sets $S_{1},S_{2}\subseteq N$ and $i\in[k]$, then by the submodularity of $f_{i}$ we can update $y_{i}(S_{1}):=y_{i}(S_{2}):=0$ and $y_{i}(S_{1}\cup S_{2}):=1$ without increasing the objective value. Hence, after such uncrossing operations, any integer feasible solution to LP-Rel induces a partition of $N$ for MSCA.
The linear program LP-Rel is equivalent to the convex program relaxation introduced by Chekuri and Ene [5] (see Appendix A for details). The linear program LP-Rel can also be seen as the dual linear program of the unweighted $k$-polymatroid intersection (see [16] and Appendix B for details). Regarding the integrality gap of LP-Rel, Ene and Vondrák [8] showed that if $\mathrm{P}\neq\mathrm{NP}$, then the integrality gap of LP-Rel for general submodular functions cannot be upper-bounded by any factor. Hence, it is natural to focus on special cases such as a monotone case that admit a bounded integrality gap. We refer to LP-Rel restricted to monotone submodular functions $f_{1},\ldots,f_{k}$ as Mono-LP-Rel. Svitkina and Tardos [18] gave a $(1+\ln n)$-approximation algorithm for Mono-MSCA, and Chekuri and Ene [5] showed that the integrality gap of Mono-LP-Rel is at most $(1+\ln n)$. Santiago and Shepherd [15] gave a $k$-approximation algorithm for Mono-MSCA and showed that the integrality gap of Mono-LP-Rel is at most $k$. On the negative side, Svitkina and Tardos [18] showed that the set cover problem with a collection of $k$ sets can be reduced to Mono-MSCA. This implies that if $k$ is a part of input, Mono-MSCA cannot be approximated within a factor of $(1-o(1))\cdot\ln n$ [10], which matches the approximation ratio shown by Svitkina and Tardos [18]. On the other hand, if $k\leq\log n$, then the set cover problem with a collection of $k$ sets can be solved in time polynomial in $n$ by brute-force. Hence, in this case the reduction does not yield any approximation lower bound for Mono-MSCA. Therefore, the approximability of Mono-MSCA and the integrality gap of Mono-LP-Rel for small $k$ are much less understood. In this paper, we aim to obtain tight upper and lower bounds of the integrality gap of Mono-LP-Rel when $k$ is small.
Our results and techniques.
As mentioned previously, the current best upper bound on the integrality gap of Mono-LP-Rel is $\min\{k,1+\ln n\}$. Our first result improves this to $k/2$.
Theorem 1.1.
The integrality gap of Mono-LP-Rel is at most $k/2$.
When $k=2$, Mono-LP-Rel coincides with the dual linear program of the unweighted polymatroid intersection. Hence, when $k=2$, Theorem 1.1 corresponds to the (total) dual integrality of the polymatroid intersection. Since the proof of Theorem 1.1 is constructive, it yields a $k/2$-approximation algorithm for Mono-MSCA, which improves the current best approximation ratio $\min\{k,1+\ln n\}$.
Theorem 1.2.
There exists a $k/2$-approximation algorithm for Mono-MSCA.
Our next result is a lower bound on the integrality gap of Mono-LP-Rel, which is tight when $k$ is fixed.
Theorem 1.3.
If $\log n\geq k$, then the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
We note that if $k$ is fixed, Theorem 1.3 implies that the integrality gap of Mono-LP-Rel is at least $k/2-\epsilon$ for any constant $\epsilon>0$, which matches the upper bound in Theorem 1.1. Another consequence is that if $k\leq\sqrt{\log n}$, then the lower bound in Theorem 1.3 satisfies $\frac{k}{2}-\frac{k(k-1)}{\log n}\geq\frac{k}{2}-1$, which almost matches the upper bound $k/2$ in Theorem 1.1.
Our proof of Theorem 1.1 is constructive as follows. We first obtain an optimal solution $(y_{1},\ldots,y_{k})$ to Mono-LP-Rel whose supports are chains via the ellipsoid method. Then, we take the top $2/k$ of a fraction of the largest sets in the chains. We show that such a collection of sets can be partitioned into groups of $k$ sets, each of which has union $N$. The existence of this partition implies that the integrality gap of Mono-LP-Rel is at most $k/2$. In the proof of Theorem 1.3, we construct $k$ monotone submodular functions such that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
Related work.
Recently, Bi et al. [3] considered the monotone submodular multiway partition problem (Mono-Sub-MP), which is a special case of Mono-MSCA. They showed that Mono-Sub-MP admits a $4/3$-approximation, and does not admit a $(10/9-\epsilon)$-approximation for every constant $\epsilon>0$. For further recent works related to submodular partition problems, see [1, 2, 4, 11, 14].
There are several recent works related to MSCA. Etesami [9] showed that multi-item welfare maximization under network externalities can be formulated as a special case of MSCA. Deng et al. [7] introduced the generalized load-balancing (GLB) problem, and showed that GLB is equivalent to MSCA when the outer norm is the $\ell_{1}$-norm.
Organization.
The rest of the paper is organized as follows. In Section 2, we show that the integrality gap of Mono-LP-Rel is at most $k/2$, which yields a $k/2$-approximation algorithm for Mono-MSCA. In Section 3, we show that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$ when $\log n\geq k$.
2 Approximation algorithm
In this section, we show that the integrality gap of Mono-LP-Rel is at most $k/2$, which yields a $k/2$-approximation algorithm for Mono-MSCA.
Let $N$ be a finite set of cardinality $n$, and $k\geq 2$ a positive integer. For each $i\in[k]$, let $f_{i}:2^{N}\to\mathbb{R}_{+}$ be a monotone submodular function with $f_{i}(\emptyset)=0$. Recall that Mono-LP-Rel is defined as follows:
$(\textsc{Mono-LP-Rel})\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
$\mathrm{subject\ to}$ $\ \ \ y_{i}\in\mathbb{R}^{2^{N}}_{+}\ \mathrm{for\ every\ }i\in[k],$
$\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)\chi_{S}=\mathbf{1}.$
To prove the upper bound of the integrality gap, we show that an integer solution to Mono-LP-Rel achieving a $k/2$-approximation can be constructed from an optimal solution to Mono-LP-Rel. Since the coefficient matrix of the constraints in Mono-LP-Rel is rational, Mono-LP-Rel has a rational optimal solution $(y^{*}_{1},\ldots,y^{*}_{k})$. Let $\mathcal{L}_{i}:=\{S\subseteq N\mid y^{*}_{i}(S)>0\}$ and $d_{i}:=|\mathcal{L}_{i}|$ for each $i\in[k]$. By the uncrossing operations, we may assume that $\mathcal{L}_{i}$ is a chain for each $i\in[k]$. Let $C_{i}^{1}\supsetneq C_{i}^{2}\supsetneq\cdots\supsetneq C_{i}^{d_{i}}$ be the members of $\mathcal{L}_{i}$ for each $i\in[k]$. Let $M$ be the product of k(k1)k(k-1) and the denominators of $y^{*}_{i}(C_{i}^{j})$ for all $i,j$. Then, $M\cdot y^{*}_{i}$ is an integer vector for every $i\in[k]$. For each $i\in[k]$ and $j\in[M]$, define $U_{i}^{j}\subseteq N$ as follows:
$U_{i}^{j}:=\begin{cases}C_{i}^{\ell}&\mathrm{if\ }1+\displaystyle\sum_{p=1}^{\ell-1}M\cdot y_{i}^{*}(C_{i}^{p})\leq j\leq\displaystyle\sum_{p=1}^{\ell}M\cdot y_{i}^{*}(C_{i}^{p})\ \mathrm{for\ some\ }\ell\in[d_{i}],\\ \emptyset&\mathrm{if\ }1+\displaystyle\sum_{p=1}^{d_{i}}M\cdot y_{i}^{*}(C_{i}^{p})\leq j\leq M.\end{cases}$
Intuitively, for each $i\in[k]$, we consider the multiset that contains $M\cdot y^{*}_{i}(C_{i}^{\ell})$ copies of $C_{i}^{\ell}$ for all $\ell\in[d_{i}]$, and $U_{i}^{j}$ is the $j$-th largest set in the multiset for each $j\in[M]$. By the definition of $U_{i}^{j}$, we have the following lemma.
Lemma 2.1.
For each $i\in[k]$ and $\ell=1,\ldots,d_{i}$, the number of $j\in[M]$ with $U_{i}^{j}=C_{i}^{\ell}$ is $M\cdot y_{i}^{*}(C_{i}^{\ell})$.
We now aim to construct an integer solution to Mono-LP-Rel (i.e., a partition of $N$) from $U_{i}^{j}$s that achieves $k/2$-approximation. To this end, we prepare the following two lemmas.
Lemma 2.2.
We have
$\sum_{i\in[k],j\in[M]}\dfrac{1}{M}f_{i}(U_{i}^{j})=\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S).$
Proof.
By Lemma 2.1, we have
$\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)=\sum_{S\in\mathcal{L}_{i},i\in[k]}y^{*}_{i}(S)f_{i}(S)=\sum_{1\leq\ell\leq d_{i},i\in[k]}y^{*}_{i}(C_{i}^{\ell})f_{i}(C_{i}^{\ell})$
$=\sum_{1\leq\ell\leq d_{i},i\in[k]}\dfrac{1}{M}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\right\}\right|\cdot f_{i}(C_{i}^{\ell})$
$=\sum_{\begin{subarray}{c}i\in[k],j\in[M]\\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }1\leq\ell\leq d_{i}\end{subarray}}\dfrac{1}{M}f_{i}(U_{i}^{j})=\sum_{i\in[k],j\in[M]}\dfrac{1}{M}f_{i}(U_{i}^{j}).$
The last equality follows since if $U_{i}^{j}\neq C_{i}^{\ell}$ for all $\ell=1,\ldots,d_{i}$, then we have $U_{i}^{j}=\emptyset$ by the definition. ∎
Lemma 2.3.
Let $a_{1},\ldots,a_{k}\in[M]$ be positive integers such that $\sum_{i=1}^{k}a_{i}\leq M+k-1$. Then, $\bigcup_{i=1}^{k}U_{i}^{a_{i}}=N$.
Proof.
Suppose to the contrary that $\bigcup_{i=1}^{k}U_{i}^{a_{i}}\neq N$. Then, there exists $e\in N$ such that $e\notin U_{i}^{a_{i}}$ for every $i\in[k]$. Combined with $U_{i}^{1}\supseteq U_{i}^{2}\supseteq\cdots\supseteq U_{i}^{M}$ for each $i\in[k]$, this implies $e\notin U_{i}^{j}$ for every $i\in[k]$ and integer $j$ with $a_{i}\leq j\leq M$. Hence, by Lemma 2.1 we have
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}M\cdot y^{*}_{i}(S)=\sum_{\begin{subarray}{c}S\in\mathcal{L}_{i},e\in S\\ i\in[k]\end{subarray}}M\cdot y^{*}_{i}(S)=\sum_{\begin{subarray}{c}1\leq\ell\leq d_{i},e\in C_{i}^{\ell}\\ i\in[k]\end{subarray}}M\cdot y^{*}_{i}(C_{i}^{\ell})$
$=\sum_{\begin{subarray}{c}1\leq\ell\leq d_{i},e\in C_{i}^{\ell}\\ i\in[k]\end{subarray}}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\right\}\right|$
$=\sum_{i\in[k]}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }\ell\in[d_{i}],\ e\in U_{i}^{j}\right\}\right|$
$\leq\sum_{i\in[k]}\left|\left\{j\in[a_{i}-1]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }\ell\in[d_{i}]\right\}\right|$
$\leq\sum_{i\in[k]}(a_{i}-1)\leq M-1.$ (2.1)
Since $y^{*}_{1},\ldots,y^{*}_{k}$ is a feasible solution to Mono-LP-Rel, we have
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}y_{i}^{*}(S)=1.$
This implies
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}M\cdot y_{i}^{*}(S)=M,$
which contradicts the inequality (2). ∎
Since $M$ is divisible by k(k1)k(k-1), we can denote $M=k(k-1)\cdot m$ for some integer $m$. For $k$ positive integers $a_{1},\ldots,a_{k}\in[M]$, we say that an ordered tuple $(a_{1},\ldots,a_{k})$ covers $N$ if $\bigcup_{i=1}^{k}U_{i}^{a_{i}}=N$. We show that if there exist distinct $M\cdot 2/k=2m(k-1)$ ordered tuples covering $N$, then one of them corresponds to an integer solution to Mono-LP-Rel that achieves $k/2$-approximation.
Lemma 2.4.
Suppose that there exist 2m(k1)2m(k-1) ordered tuples
$(a_{1}^{1},\ldots,a_{k}^{1}),(a_{1}^{2},\ldots,a_{k}^{2}),\ldots,(a_{1}^{2m(k-1)},\ldots,a_{k}^{2m(k-1)})$
such that
  • (1)
    $a_{i}^{j}\in[M]$ for each $i\in[k]$ and $j\in[2m(k-1)]$,
  • (2)
    $a_{i}^{j}\neq a_{i}^{\ell}$ for each $i\in[k]$ and $j,\ell\in[2m(k-1)]$ with $j\neq\ell$,
  • (3)
    each ordered tuple covers $N$.
Then, the integrality gap of Mono-LP-Rel is at most $k/2$.
Proof.
By Lemma 2.2, we have
$\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)=\sum_{i\in[k],j\in[M]}\dfrac{1}{M}f_{i}(U_{i}^{j})\geq\sum_{i\in[k],j\in[2m(k-1)]}\dfrac{1}{M}f_{i}(U_{i}^{a_{i}^{j}})$
$\geq\dfrac{2m(k-1)}{M}\cdot\min\left\{\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}})\ \middle|\ j\in[2m(k-1)]\right\}$
$=\dfrac{2}{k}\cdot\min\left\{\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}})\ \middle|\ j\in[2m(k-1)]\right\}.$ (2.2)
The first inequality follows since by the monotonicity of $f_{i}$, we have $f_{i}(U_{i}^{j})\geq f_{i}(\emptyset)=0$ for every $i\in[k]$ and $j\in[M]$. Let
$j^{*}:=\underset{j\in[2m(k-1)]}{\operatorname{argmin}}\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}}).$
Then, by (2) we have
$\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\geq\dfrac{2}{k}\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j^{*}}}).$
Hence, we have
$\dfrac{k}{2}\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\geq\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j^{*}}}).$ (2.3)
Since $\bigcup_{i=1}^{k}U_{i}^{a_{i}^{j^{*}}}=N$, there exists a partition $U_{1},\ldots,U_{k}$ of $N$ such that $U_{i}\subseteq U_{i}^{a_{i}^{j^{*}}}$ for every $i\in[k]$. Then, by (2.3) and the monotonicity of $f_{i}$, we have
$\dfrac{k}{2}\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\geq\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j^{*}}})\geq\sum_{i\in[k]}f_{i}(U_{i}).$
This implies that the integrality gap of Mono-LP-Rel is at most $k/2$. ∎
We are now ready to prove Theorem 1.1.
Theorem 1.1.
By Lemma 2.4, it suffices to show that there exist 2m(k1)2m(k-1) ordered tuples satisfying the conditions in Lemma 2.4. For each positive integer $a\in\mathbb{Z}_{>0}$, define $\mathrm{mod}(a)$ as follows:
[…]
For each $i\in[k]$ and $j\in[2m(k-1)]$, define $a_{i}^{j}$ as follows:
$a_{i}^{j}:=\begin{cases}\mathrm{mod}(2m(i-1)+j)&\mathrm{if\ }1\leq i\leq k-1,\\ \mathrm{mod}((2m-j+1)(k-1))-\left\lfloor\dfrac{j-1}{2m}\right\rfloor&\mathrm{if\ }i=k.\end{cases}$
We show that $(a_{i}^{j})_{i\in[k],j\in[2m(k-1)]}$ satisfies conditions (1)–(3) in Lemma 2.4. For condition (1), we show that $1\leq a_{i}^{j}\leq M$ for each $i\in[k]$ and $j\in[2m(k-1)]$. If $1\leq i\leq k-1$, then we have
$1\leq a_{i}^{j}=\mathrm{mod}(2m(i-1)+j)\leq 2m(k-1)\leq km(k-1)=M.$
If $i=k$, then we have
$a_{i}^{j}=\mathrm{mod}((2m-j+1)(k-1))-\left\lfloor\dfrac{j-1}{2m}\right\rfloor.$
Since (2mj+1)(k1)(2m-j+1)(k-1) is a multiple of k1k-1, we have
$\mathrm{mod}((2m-j+1)(k-1))=p(k-1)$
for some positive integer $p$. This implies
$\mathrm{mod}((2m-j+1)(k-1))\geq k-1.$
Hence, we have
$1=k-1-(k-2)=k-1-\left\lfloor\dfrac{2m(k-1)-1}{2m}\right\rfloor\leq\mathrm{mod}((2m-j+1)(k-1))-\left\lfloor\dfrac{j-1}{2m}\right\rfloor,$
and
$\mathrm{mod}((2m-j+1)(k-1))-\left\lfloor\dfrac{j-1}{2m}\right\rfloor\leq 2m(k-1)\leq km(k-1)=M.$
Therefore, condition (1) holds for $(a_{i}^{j})_{i\in[k],j\in[2m(k-1)]}$.
For condition (2), we next show that $a_{i}^{j}\neq a_{i}^{\ell}$ for each $i\in[k]$ and $j,\ell\in[2m(k-1)]$ with $j\neq\ell$. If $1\leq i\leq k-1$, then by the definition of $a_{i}^{j}$, it is clear that $a_{i}^{j}\neq a_{i}^{\ell}$ holds for each $j,\ell\in[2m(k-1)]$ with $j\neq\ell$. Consider the case when $i=k$. Suppose to the contrary that $a_{k}^{j}=a_{k}^{\ell}$ holds for some $j,\ell\in[2m(k-1)]$ with $j\neq\ell$. This implies that
$\mathrm{mod}((2m-j+1)(k-1))-\left\lfloor\dfrac{j-1}{2m}\right\rfloor=\mathrm{mod}((2m-\ell+1)(k-1))-\left\lfloor\dfrac{\ell-1}{2m}\right\rfloor.$ (2.4)
Since $\mathrm{mod}((2m-j+1)(k-1))$ and $\mathrm{mod}((2m-\ell+1)(k-1))$ are multiples of k1k-1, it follows that
$\left\lfloor\dfrac{j-1}{2m}\right\rfloor-\left\lfloor\dfrac{\ell-1}{2m}\right\rfloor$
is a multiple of k1k-1 by (2.4). Combined with
$0\leq\left\lfloor\dfrac{j-1}{2m}\right\rfloor,\left\lfloor\dfrac{\ell-1}{2m}\right\rfloor\leq k-2,$
this implies that
$\left\lfloor\dfrac{j-1}{2m}\right\rfloor=\left\lfloor\dfrac{\ell-1}{2m}\right\rfloor.$
Hence, we have $2m(p-1)+1\leq j,\ell\leq 2mp$ for some $p\in[k-1]$. Let $j^{\prime}:=2mp-j$ and $\ell^{\prime}:=2mp$-$\ell$. By the definition of $p$, we have $0\leq j^{\prime},\ell^{\prime}\leq 2m-1$. By (2.4), we have
$\mathrm{mod}((2m-j+1)(k-1))=\mathrm{mod}((2m-\ell+1)(k-1)).$ (2.5)
We also have
$\mathrm{mod}((2m-j+1)(k-1))=\mathrm{mod}((2m-2mp+j^{\prime}+1)(k-1))=(j^{\prime}+1)(k-1),$
$\mathrm{mod}((2m-\ell+1)(k-1))=\mathrm{mod}((2m-2mp+\ell^{\prime}+1)(k-1))=(\ell^{\prime}+1)(k-1).$ (2.6)
Hence, by (2.5) and (2) we have
$(j^{\prime}+1)(k-1)=(\ell^{\prime}+1)(k-1).$
This implies $j^{\prime}=\ell^{\prime}$, which contradicts $j\neq\ell$.
For condition (3), we show that $\bigcup_{i=1}^{k}U_{i}^{a_{i}^{j}}=N$ holds for every $j\in[2m(k-1)]$. By Lemma 2.3, it suffices to show that $\sum_{i=1}^{k}a_{i}^{j}\leq M+k-1$ for every $j\in[2m(k-1)]$. Take any $j\in[2m(k-1)]$. Let $p\in[k-1]$ be the unique positive integer such that $2m(p-1)+1\leq j\leq 2mp$. Let $j^{\prime}:=2mp-j$. Then, we have $0\leq j^{\prime}\leq 2m-1$. By the definition of $a_{i}^{j}$, we have
$\sum_{i=1}^{k}a_{i}^{j}=\sum_{i=1}^{k-1}\mathrm{mod}(2m(i-1)+j)+\mathrm{mod}((2m-j+1)(k-1))-\left\lfloor\dfrac{j-1}{2m}\right\rfloor$
$=\sum_{i=1}^{k-1}\mathrm{mod}(2m(i-1)+2mp-j^{\prime})+\mathrm{mod}((2m-2mp+j^{\prime}+1)(k-1))-\left\lfloor\dfrac{2mp-j^{\prime}-1}{2m}\right\rfloor$
$=\sum_{i=1}^{k-p}(2m(i-1)+2mp-j^{\prime})+\sum_{i=k-p+1}^{k-1}(2m(i-1)+2mp-j^{\prime}-2m(k-1))$
$+(j^{\prime}+1)(k-1)-(p-1)$
$=\sum_{i=1}^{k-1}(2m(i-1)+2mp-j^{\prime})-2m(k-1)(p-1)+(j^{\prime}+1)(k-1)-(p-1)$
$=mk(k-1)-(k-1)(2m-2mp+j^{\prime})-2m(k-1)(p-1)+(j^{\prime}+1)(k-1)-(p-1)$
$=M+k-p\leq M+k-1.$
This completes the proof. ∎
The proof of Theorem 1.1 provides the construction of a $k/2$-approximate integer solution to Mono-LP-Rel. We prove Theorem 1.2 by showing that such an integer solution can be constructed in polynomial time.
Theorem 1.2.
We can obtain a rational optimal solution $(y_{1}^{*},\ldots,y_{k}^{*})$ to Mono-LP-Rel whose supports $\mathcal{L}_{1},\ldots,\mathcal{L}_{k}$ are chains in polynomial time via the ellipsoid method [5] (see Appendix A for details). Then, we define $U_{i}^{j}$ for each $i\in[k]$ and $j\in[M]$, and $a_{i}^{j}$ for each $i\in[k]$ and $j\in[2m(k-1)]$ as in the proof of Theorem 1.1. If we can compute
$j^{*}:=\underset{j\in[2m(k-1)]}{\operatorname{argmin}}\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}}),$
then any partition $U_{1},\ldots,U_{k}$ of $N$ such that $U_{i}\subseteq U_{i}^{a_{i}^{j^{*}}}$ for every $i\in[k]$ satisfies
$\sum_{i\in[k]}f_{i}(U_{i})\leq\dfrac{k}{2}\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\leq\dfrac{k}{2}\cdot\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\},$
as desired. Hence, it suffices to show that $j^{*}$ can be computed in polynomial time. To find $j^{*}$, we set $j_{1}:=1$. Then, for q=2,3,,q=2,3,..., we define $j_{q}$ as the minimum integer such that $j_{q}>j_{q-1}$ and there exists $i\in[k]$ with $U_{i}^{a_{i}^{j_{q}}}\neq U_{i}^{a_{i}^{j_{q-1}}}$. Recall that $n:=|N|$. Since for each $i\in[k]$, we have $U_{i}^{1}\supseteq U_{i}^{2}\supseteq\cdots\supseteq U_{i}^{M}$ and there are at most $|\mathcal{L}_{i}|\leq n$ distinct sets among $\{U_{i}^{j}\}_{j=1}^{M}$, we only need to define $j_{q}$ for $q=1,2,\ldots,Q$ with $Q\leq kn$. Since we have
$j^{*}=\underset{j\in\{j_{1},j_{2},\ldots,j_{Q}\}}{\operatorname{argmin}}\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}}),$
we can find $j^{*}$ by computing the values of $\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j_{q}}})$ for all $q\in[Q]$. By the definition of $U_{i}^{j}$ and $a_{i}^{j}$, we can compute $j_{q}$ in polynomial time for each fixed $q$. Hence, since $Q\leq kn$, we can compute $j_{q}$ and $\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j_{q}}})$ for all $q\in[Q]$ in polynomial time. ∎
3 Lower bound on the integrality gap
In this section, we show that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$ when $\log n\geq k$. To this end, we explicitly construct monotone submodular functions $f_{1},\ldots,f_{k}$ such that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
Suppose that $\log n\geq k$. Let $p$ be the unique positive integer such that $pk\leq\lfloor\log n\rfloor\leq(p+1)k-1$. We define
$\mathcal{V}:=\left\{v\in\mathbb{Z}_{+}^{k}\ \middle|\ \sum_{i=1}^{k}v_{i}=pk-k+1\right\}.$
By stars and bars (see e.g., [17, Chapter 1]), we have
$|\mathcal{V}|=\binom{pk}{k-1}\leq 2^{pk}\leq 2^{\lfloor\log n\rfloor}\leq n=|N|.$
Choose $e_{v}\in N$ for each $v\in\mathcal{V}$ so that all $e_{v}$ are distinct. For each $i\in[k]$, let
$N_{i}:=\{e_{v}\in N\mid v\in\mathcal{V},v_{i}\geq 1\},$
$Z_{i}:=\{e_{v}\in N\mid v\in\mathcal{V},v_{i}=0\}.$
For each $i\in[k]$, define $f_{i}:2^{N}\to\mathbb{R}$ as follows: for all $S\subseteq N$,
$f_{i}(S):=\begin{cases}\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in S\cap N_{i}\}\}&\mathrm{if\ }S\cap N_{i}\neq\emptyset\ \mathrm{and\ }S\cap Z_{i}=\emptyset,\\ 0&\mathrm{if\ }S\cap N_{i}=\emptyset\ \mathrm{and\ }S\cap Z_{i}=\emptyset,\\ 2pk+1&\mathrm{if\ }S\cap Z_{i}\neq\emptyset.\end{cases}$
Lemma 3.1.
$f_{i}$ is monotone submodular for each $i\in[k]$.
Proof.
We first show monotonicity of $f_{i}$. Take any $S,T\subseteq N$ with $S\subseteq T$. If $S\cap N_{i}=S\cap Z_{i}=\emptyset$, then we have $f_{i}(S)=0\leq f_{i}(T)$. If $S\cap Z_{i}\neq\emptyset$, then we have $T\cap Z_{i}\neq\emptyset$, which implies $f_{i}(S)=f_{i}(T)=2pk+1$. Consider the case when $S\cap N_{i}\neq\emptyset$ and $S\cap Z_{i}=\emptyset$. If $T\cap Z_{i}\neq\emptyset$, then we have
$f_{i}(S)=\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in S\cap N_{i}\}\}\leq 2p+1-1=2p\leq 2pk+1=f_{i}(T).$
If $T\cap Z_{i}=\emptyset$, then since $T\cap N_{i}\neq\emptyset$ follows from $S\cap N_{i}\neq\emptyset$, we have
$f_{i}(S)$ $=\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in S\cap N_{i}\}\}$
$\leq\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in T\cap N_{i}\}\}=f_{i}(T).$
Therefore, $f_{i}$ is monotone.
We next show that $f_{i}$ is submodular. Take any $S,T\subseteq N$. If $S\cap Z_{i}\neq\emptyset$, then by the monotonicity of $f_{i}$, we have
$f_{i}(S)+f_{i}(T)=2pk+1+f_{i}(T)\geq 2pk+1+f_{i}(S\cap T)\geq f_{i}(S\cup T)+f_{i}(S\cap T).$
The last inequality follows since $f_{i}(X)\leq 2pk+1$ holds for every $X\subseteq N$ by the definition of $f_{i}$. If $S\cap N_{i}=S\cap Z_{i}=\emptyset$, then since $(S\cup T)\cap N_{i}=T\cap N_{i}$ and $(S\cup T)\cap Z_{i}=T\cap Z_{i}$, we have $f_{i}(S\cup T)=f_{i}(T)$ by the definition of $f_{i}$. This implies that
$f_{i}(S)+f_{i}(T)=f_{i}(S)+f_{i}(S\cup T)\geq f_{i}(S\cup T)+f_{i}(S\cap T).$
The last inequality follows from the monotonicity of $f_{i}$. Hence, by the symmetry between $S$ and $T$, the remaining case is when $S\cap N_{i}\neq\emptyset$ and $S\cap Z_{i}=\emptyset$, and $T\cap N_{i}\neq\emptyset$ and $T\cap Z_{i}=\emptyset$. In this case, we have $(S\cup T)\cap N_{i}\neq\emptyset$ and $(S\cup T)\cap Z_{i}=\emptyset$, which implies that
$f_{i}(S\cup T)=\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in(S\cup T)\cap N_{i}\}\}.$
Take $e_{v^{*}}\in\mathrm{argmax}\{2p+1-v_{i}\mid e_{v}\in(S\cup T)\cap N_{i}\}$ with the corresponding vector $v^{*}\in\mathcal{V}$. Then, we have $e_{v^{*}}\in S\cap N_{i}$ or $e_{v^{*}}\in T\cap N_{i}$. We may assume without loss of generality that $e_{v^{*}}\in S\cap N_{i}$. Then, we have $e_{v^{*}}\in\mathrm{argmax}\{2p+1-v_{i}\mid e_{v}\in S\cap N_{i}\}$. Hence, we have
$f_{i}(S)+f_{i}(T)=\max\{0,2p+1-v^{*}_{i}\}+f_{i}(T)=f_{i}(S\cup T)+f_{i}(T)$ $\geq f_{i}(S\cup T)+f_{i}(S\cap T).$
The last inequality follows from the monotonicity of $f_{i}$. Therefore, $f_{i}$ is submodular. ∎
We now show that for the monotone submodular functions $f_{1},\ldots,f_{k}$ defined above, any optimal solution $(y_{1}^{*},\ldots,y_{k}^{*})$ to Mono-LP-Rel satisfies
$\left(\frac{k}{2}-\frac{k(k-1)}{\log n}\right)\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)$
$\leq\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$ (3.1)
The inequality (3) implies that the integrality gap of Mono-LP-Rel is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$, which completes the proof of Theorem 1.3.
To prove (3), we first give an upper bound on the value $\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)$. To this end, we construct a feasible solution to Mono-LP-Rel. For each $i\in[k]$ and $j\in[pk-k+1]$, let $N_{i}^{j}:=\{e_{v}\in N\mid v\in\mathcal{V},v_{i}\geq j\}$. Note that $N_{i}^{1}=N_{i}$. In what follows, we will use $N_{i}^{1}$ instead of $N_{i}$ for convenience. Let $N^{c}:=N\setminus\{e_{v}\in N\mid v\in\mathcal{V}\}$. For each $i\in[k]$, define $z_{i}\in\mathbb{R}_{+}^{2^{N}}$ as follows: for all $S\subseteq N$,
$z_{i}(S):=\begin{cases}\dfrac{1}{pk-k+1}&\mathrm{if\ }i=1\ \mathrm{and\ }\vskip 3.0ptS=N_{1}^{j}\cup N^{c}\ \mathrm{for\ some\ }j\in[pk-k+1],\\ \vskip 3.0pt\dfrac{1}{pk-k+1}&\mathrm{if\ }2\leq i\leq k\ \mathrm{and\ }S=N_{i}^{j}\ \mathrm{for\ some\ }j\in[pk-k+1],\\ 0&\mathrm{otherwise}.\end{cases}$
Lemma 3.2.
$(z_{1},\ldots,z_{k})$ is a feasible solution to Mono-LP-Rel.
Proof.
To show the feasibility of $(z_{1},\ldots,z_{k})$, we show that
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}z_{i}(S)=1$ (3.2)
for every $e\in N$. For any $e\in N^{c}$, we have
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}z_{i}(S)=\sum_{j\in[pk-k+1]}z_{1}(N_{1}^{j}\cup N^{c})=\sum_{j\in[pk-k+1]}\dfrac{1}{pk-k+1}=1.$
Take any $e\in N\setminus N^{c}$. Since $e=e_{v}$ for some $v\in\mathcal{V}$, we have
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}z_{i}(S)$ $=\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ v_{1}\geq j\end{subarray}}z_{1}(N_{1}^{j}\cup N^{c})+\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ 2\leq i\leq k,v_{i}\geq j\end{subarray}}z_{i}(N_{i}^{j})$
$=\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ v_{1}\geq j\end{subarray}}\frac{1}{pk-k+1}+\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ 2\leq i\leq k,v_{i}\geq j\end{subarray}}\frac{1}{pk-k+1}$
$=\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ i\in[k],v_{i}\geq j\end{subarray}}\frac{1}{pk-k+1}$
$=\frac{pk-k+1}{pk-k+1}=1.$
The second last equality follows since $\sum_{i=1}^{k}v_{i}=pk-k+1$. Hence, (3.2) holds for every $e\in N$, which implies that $(z_{1},\ldots,z_{k})$ is a feasible solution to Mono-LP-Rel. ∎
By Lemma 3.2 and the optimality of $y^{*}_{1},\ldots,y^{*}_{k}$, we have
$\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\leq\sum_{S\subseteq N,i\in[k]}z_{i}(S)f_{i}(S)$
$=\sum_{j\in[pk-k+1]}\frac{1}{pk-k+1}f_{1}(N_{1}^{j}\cup N^{c})+\sum_{2\leq i\leq k,j\in[pk-k+1]}\frac{1}{pk-k+1}f_{i}(N_{i}^{j})$
$=\sum_{j\in[pk-k+1]}\frac{1}{pk-k+1}\cdot\max\{0,2p+1-j\}+\sum_{2\leq i\leq k,j\in[pk-k+1]}\frac{1}{pk-k+1}\cdot\max\{0,2p+1-j\}$
$=\frac{pk(2p+1)}{pk-k+1}.$ (3.3)
To prove (3), we next provide a lower bound on the right hand side of (3).
Lemma 3.3.
Let $X_{1},\ldots,X_{k}$ be any partition of $N$. Then, we have
$\sum_{i=1}^{k}f_{i}(X_{i})\geq pk+k.$
Proof.
If $X_{j}\cap Z_{j}\neq\emptyset$ for some $j\in[k]$, then we have
$\sum_{i=1}^{k}f_{i}(X_{i})\geq f_{j}(X_{j})=2pk+1\geq pk+k.$
Hence, it suffices to consider the case when $X_{i}\cap Z_{i}=\emptyset$ for every $i\in[k]$. For each $i\in[k]$, define $x^{i}\in\mathcal{V}$ as follows:
$x^{i}_{j}:=\begin{cases}pk-k+1&\mathrm{if\ }j=i,\\ 0&\mathrm{if\ }j\in[k]\mathrm{\ and\ }j\neq i.\end{cases}$
Then, since $X_{i}\cap Z_{i}=\emptyset$ for each $i\in[k]$, we have $e_{x^{i}}\in X_{i}$ for every $i\in[k]$. Hence, we have $X_{i}\cap N_{i}\neq\emptyset$ for every $i\in[k]$. For each $i\in[k]$, we define $b_{i}:=\min\{v_{i}\mid e_{v}\in X_{i}\cap N_{i}\}$. We now show that $\sum_{i=1}^{k}b_{i}\leq pk$. Suppose to the contrary that $\sum_{i=1}^{k}b_{i}\geq pk+1$. Then, since $\sum_{i=1}^{k}(b_{i}-1)\geq pk-k+1$, there exists $v\in\mathcal{V}$ such that $v_{i}\leq b_{i}-1$ for every $i\in[k]$. Choose $i\in[k]$ such that $e_{v}\in X_{i}$. Since $X_{i}\cap Z_{i}=\emptyset$ by the assumption, we have $e_{v}\in X_{i}\cap N_{i}$, which contradicts the definition of $b_{i}$. Hence, we have $\sum_{i=1}^{k}b_{i}\leq pk$. Therefore, since $X_{i}\cap N_{i}\neq\emptyset$ and $X_{i}\cap Z_{i}=\emptyset$ for each $i\in[k]$, we have
$\sum_{i=1}^{k}f_{i}(X_{i})=\sum_{i=1}^{k}\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in X_{i}\cap N_{i}\}\}$
$\geq\sum_{i=1}^{k}\max\{2p+1-v_{i}\mid e_{v}\in X_{i}\cap N_{i}\}=2pk+k-\sum_{i=1}^{k}\min\{v_{i}\mid e_{v}\in X_{i}\cap N_{i}\}$
$=2pk+k-\sum_{i=1}^{k}b_{i}\geq 2pk+k-pk=pk+k.$
By Lemma 3.3, we have
$\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}$
$\geq pk+k=\frac{2k(p+1)^{2}(pk-k+1)}{2(p+1)(pk-k+1)}$
$=\frac{2k(p+1)(p^{2}k-k+p+1)}{2(p+1)(pk-k+1)}=\frac{2k(p+1)(p^{2}k+(p-1)(k-1)-pk+2p)}{2(p+1)(pk-k+1)}$
$\geq\frac{2k(p+1)(p^{2}k-pk+2p)}{2(p+1)(pk-k+1)}=\frac{pk(2p+2)(pk-k+2)}{2(p+1)(pk-k+1)}$
$\geq\frac{pk(2p+1)(pk-k+2)}{2(p+1)(pk-k+1)}=\frac{pk(2p+1)}{pk-k+1}\cdot\left(\frac{k}{2}-\frac{k(k-1)}{k(p+1)}\right)$
$\geq\frac{pk(2p+1)}{pk-k+1}\cdot\left(\frac{k}{2}-\frac{k(k-1)}{\log n}\right).$ (3.4)
The last inequality follows from the definition of $p$. Combining (3) and (3), we have
$\left(\frac{k}{2}-\frac{k(k-1)}{\log n}\right)\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\leq\left(\frac{k}{2}-\frac{k(k-1)}{\log n}\right)\cdot\frac{pk(2p+1)}{pk-k+1}$
$\leq\min\left\{\sum_{i=1}^{k}f_{i}(X_{i})\ \middle|\ X_{1},\ldots,X_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$
Thus, the inequality (3) holds, which completes the proof of Theorem 1.3.
Acknowledgements
This work was supported by JST ERATO Grant Number JPMJER2301, Japan.
Appendix A Convex program based on the Lovász extension
In this section, we show that LP-Rel is equivalent to the convex program relaxation of MSCA introduced by Chekuri and Ene [5]. Let $N:=\{1,\ldots,n\}$. Let $f_{i}:2^{N}\to\mathbb{R}$ be a submodular function with $f_{i}(\emptyset)=0$ for each $i\in[k]$. Recall that LP-Rel is defined as follows:
$(\textsc{LP-Rel})\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
$\mathrm{subject\ to}$ $\ \ \ y_{i}\in\mathbb{R}^{2^{N}}_{+}\ \mathrm{for\ every\ }i\in[k],$
$\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)\chi_{S}=\mathbf{1}.$
For a set function $f:2^{N}\to\mathbb{R}$ with $f(\emptyset)=0$, the Lovász extension $\hat{f}:[0,1]^{n}\to\mathbb{R}$ of $f$ is defined as follows: Take any $x\in[0,1]^{n}$. Relabel the elements in $N$ so that $x_{1}\geq x_{2}\geq\cdots\geq x_{n}$. For convenience, let $x_{n+1}:=0$. Let $S_{i}:=\{1,2,\ldots,i\}$ for each $i\in[n]$. The value of $\hat{f}(x)$ is defined as follows:
$\hat{f}(x):=\sum_{i=1}^{n}(x_{i}-x_{i+1})f(S_{i}).$
Chekuri and Ene [5] introduced the following convex program relaxation of MSCA based on the Lovász extension:
$(\textsc{LE-Rel})\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{i=1}^{k}\hat{f}_{i}(x^{i})$
$\mathrm{subject\ to}$ $\ \ \ x^{i}\in[0,1]^{n}\ \mathrm{for\ every\ }i\in[k],$
$\ \ \ \sum_{i=1}^{k}x^{i}_{j}=1\ \mathrm{for\ every\ }j\in[n].$
We now show that any feasible solution $({x^{1}},\ldots,x^{k})$ to LE-Rel can be translated into a feasible solution $(y_{1},\ldots,y_{k})$ to LP-Rel that has the same objective value and whose supports are chains. Take any feasible solution $({x^{1}},\ldots,x^{k})$ to LE-Rel. We will describe a translation from $x^{i}$ to $y_{i}$ for each $i\in[k]$. We first relabel the elements in $N$ so that $x^{i}_{1}\geq x^{i}_{2}\geq\cdots\geq x^{i}_{n}$. Let $x^{i}_{n+1}:=0$ and $S_{j}:=\{1,2,\ldots,j\}$ for each $j\in[n]$. Then, for all $S\subseteq N$ define $y_{i}(S)$ as follows:
$y_{i}(S):=\begin{cases}x_{j}^{i}-x_{j+1}^{i}&\mathrm{if\ }S=S_{j}\ \mathrm{and\ }x_{j}^{i}-x_{j+1}^{i}>0\ \mathrm{for\ some\ }j\in[n],\\ 0&\mathrm{otherwise}.\end{cases}$
Then, $(y_{1},\ldots,y_{k})$ is a feasible solution to LP-Rel because for each $i\in[k]$ we have
$\sum_{S\subseteq N}y_{i}(S)\chi_{S}=\sum_{j\in[n]}(x^{i}_{j}-x^{i}_{j+1})\chi_{S_{j}}=x^{i}$
and
$\sum_{S\subseteq N,i\in[k]}y_{i}(S)\chi_{S}=\sum_{i\in[k]}x^{i}=\mathbf{1}.$
For each $i\in[k]$, we also have
$\sum_{S\subseteq N}y_{i}(S)f_{i}(S)=\sum_{j\in[n]}(x_{j}^{i}-x_{j+1}^{i})f_{i}(S_{j})=\hat{f_{i}}(x^{i}).$
Hence, we obtain
$\sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)=\sum_{i=1}^{k}\hat{f_{i}}(x^{i}),$
which implies that $(y_{1},\ldots,y_{k})$ has the same objective value as $(x^{1},\ldots,x^{k})$. Similarly, we can translate any feasible solution to LP-Rel whose supports are chains into a feasible solution to LE-Rel that has the same objective value.
Since LE-Rel can be solved in polynomial time via the ellipsoid method [5], an optimal solution to LP-Rel whose supports are chains can be obtained in polynomial time. Chekuri and Ene [6] provided another equivalent linear programming formulation of MSCA called LE-Rel-Primal. We note that LP-Rel is equivalent to LE-Rel-Primal.
Appendix B $k$-polymatroid intersection
Let $N$ be a finite set and $k\geq 2$ a positive integer. For each $i\in[k]$, let $f_{i}:2^{N}\to\mathbb{R}$ be a submodular function with $f_{i}(\emptyset)=0$. For a set function $f:2^{N}\to\mathbb{R}$, we define $P(f):=\{x\in\mathbb{R}^{N}\mid x(S)\leq f(S)\ \mathrm{for\ every\ }S\subseteq N\}$, where $x(S):=\sum_{e\in S}x(e)$ for every $S\subseteq N$. Let $w\in\mathbb{R}_{+}^{N}$ be a weight vector. The weighted $k$-polymatroid intersection problem is defined as follows:
$(\mathrm{LP}_{k\text{-poly}})\ \ \ \ \ \mathrm{maximize}$ $\ \ \ w^{\top}x$
$\mathrm{subject\ to}$ $\ \ \ x\in P(f_{i})\ \mathrm{for\ every\ }i\in[k].$
The linear program $\mathrm{LP}_{k\text{-poly}}$ corresponds to the weighted polymatroid intersection when $k=2$. When $w=\mathbf{1}$, the linear program $\mathrm{LP}_{k\text{-poly}}$ is the dual linear program of LP-Rel. For the case when $f_{1},\ldots,f_{k}$ are matroid rank functions (i.e., the case of weighted $k$-matroid intersection), the integrality gap of $\mathrm{LP}_{k\text{-poly}}$ is $2$ when $k=3$ [13], and it equals k1k-1 for arbitrary $k$ in the unweighted case $w=\mathbf{1}$ [12].