### v2

# An Approximation Algorithm for Monotone Submodular Cost Allocation

Abstract
In this paper, we consider the minimum submodular cost allocation (MSCA) problem. The input of MSCA is $k$ non-negative submodular functions $f_{1},f_{2},\ldots,f_{k}$ on the ground set $N$ given by evaluation oracles, and the goal is to partition $N$ into $k$ (possibly empty) sets $S_{1},S_{2},\ldots,S_{k}$ so that $\sum_{i=1}^{k}f_{i}(S_{i})$ is minimized. In this paper, we focus on the case when $f_{1},f_{2},\ldots,f_{k}$ are monotone, which coincides with the submodular facility location problem considered by Svitkina and Tardos. We show that the integrality gap of a natural LP-relaxation for MSCA with monotone submodular functions is at most $k/2$, yielding a $k/2$-approximation algorithm. We also prove a nearly matching lower bound: the integrality gap is at least $k/2-\epsilon$ for any constant $\epsilon>0$ when $k$ is fixed.
1 Introduction
Let $N$ be a finite set and $k\geq 2$ a positive integer. A set function $f:2^{N}\to\mathbb{R}$ is called submodular if $f(S)+f(T)\geq f(S\cup T)+f(S\cap T)$ holds for every $S,T\subseteq N$. In this paper, we focus on the Minimum Submodular Cost Allocation (MSCA) problem, introduced by Chekuri and Ene [5]. The input of MSCA is $k$ non-negative submodular functions $f_{1},f_{2},\ldots,f_{k}:2^{N}\to\mathbb{R}_{+}$ given by evaluation oracles, and the goal is to partition $N$ into $k$ (possibly empty) sets $S_{1},S_{2},\ldots,S_{k}$ so that $\sum_{i=1}^{k}f_{i}(S_{i})$ is minimized, i.e.,
$\min\left\{\sum_{i=1}^{k}f_{i}(S_{i})\ \middle|\ S_{1},S_{2},\ldots,S_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$
Several well-studied problems such as the multiway cut problems in graphs and hypergraphs, the submodular multiway partition problems, and uniform metric labeling can be cast as special cases of MSCA [5]. When all submodular functions are monotone, MSCA is equivalent to the submodular facility location problem considered by Svitkina and Tardos [18]. In this paper, we focus on this setting, that is, we assume that $f_{i}(S)\leq f_{i}(T)$ for all $S\subseteq T\subseteq N$ and $i=1,2,\ldots,k$. We call this monotone case of MSCA the Minimum Monotone Submodular Cost Allocation (MMSCA) problem. The aim of this paper is to investigate the integrality gap of a natural LP-relaxation for MMSCA when $k$ is small.
LP-relaxation.
To provide an LP-relaxation for MMSCA, we introduce some notation. Let $\mathbf{1}\in\mathbb{Z}^{N}_{>0}$ denote the all-ones vector and $k]:=\{1,2,\ldots,k\}$. For any $S\subseteq N$, let $\chi_{S}\in\{0,1\}^{N}$ denote the characteristic vector of $S$. Throughout this paper, we assume that the input functions $f_{1},f_{2},\ldots,f_{k}$ satisfy $f_{i}(\emptyset)=0$ for each $i=1,2,\ldots,k$. We consider the following linear programming relaxation for MSCA:
$(\textsc{LP})\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
$\mathrm{subject\ to}$ $\ \ \ y_{i}\in\mathbb{R}^{2^{N}}_{+}\ \mathrm{for\ every\ }i\in[k],$
$\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)\chi_{S}=\mathbf{1}.$
In this LP-relaxation, we interpret $y_{i}(S)=1$ to mean that $S$ is assigned to $f_{i}$ in MSCA, whereas $y_{i}(S)=0$ means that $S$ is not assigned to $f_{i}$. We note that if an integer feasible solution $(y_{1},y_{2},\ldots,y_{k})$ satisfies $y_{i}(S)=y_{i}(T)=1$ for two distinct sets $S,T\subseteq N$ and $i\in[k]$, then by the submodularity of $f_{i}$ we can update $y_{i}(S):=y_{i}(T):=0$ and $y_{i}(S\cup T):=1$ without increasing the objective value. Hence, after such uncrossing operations, any integer feasible solution to (LP) induces a partition of $N$ for MSCA.
The linear program (LP) is equivalent to the convex program relaxation introduced by Chekuri and Ene [5]. The linear program (LP) can also be seen as the dual linear program of the unweighted $k$-polymatroid intersection; see, e.g., [16]. Regarding the integrality gap of (LP), Ene and Vondrák [8] showed that if $\mathrm{P}\neq\mathrm{NP}$, then the integrality gap of (LP) for general submodular functions cannot be upper-bounded by any factor. Hence, it is natural to focus on special cases such as a monotone case that admit a bounded integrality gap. We refer to (LP) restricted to monotone submodular functions $f_{1},f_{2},\ldots,f_{k}$ as $\mathrm{(LP_{mono})}$. Svitkina and Tardos [18] gave a $(1+\ln n)$-approximation algorithm for MMSCA, and Chekuri and Ene [5] showed that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $(1+\ln n)$, where $n:=|N|$. Santiago and Shepherd [15] showed that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k$, yielding a $k$-approximation for MMSCA. On the negative side, Svitkina and Tardos [18] showed that the set cover problem with a collection of $k$ sets can be reduced to MMSCA. This implies that if $k$ is a part of input, MMSCA cannot be approximated within a factor of $(1-o(1))\cdot\ln n$ [10], which matches the approximation ratio shown by Svitkina and Tardos [18]. On the other hand, if $k\leq\log n$, then the set cover problem with a collection of $k$ sets can be solved in time polynomial in $n$ by brute-force. Hence, in this case the reduction does not yield any approximation lower bound for MMSCA. Therefore, the approximability of MMSCA and the integrality gap of $\mathrm{(LP_{mono})}$ for small $k$ are much less understood. In this paper, we aim to obtain tight upper and lower bounds of the integrality gap of $\mathrm{(LP_{mono})}$ when $k$ is small.
Our results and techniques.
As mentioned previously, the current best upper bound on the integrality gap of $\mathrm{(LP_{mono})}$ is $\min\{k,1+\ln n\}$. Our first result improves this to $k/2$.
Theorem 1.1.
The integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$.
Recall that when $k=2$, $\mathrm{(LP_{mono})}$ is the dual linear program of the unweighted polymatroid intersection. Since the constraint system of the unweighted polymatroid intersection is totally dual integral (see, e.g., [16]), its dual admits an integral optimal solution, which matches Theorem 1.1 for $k=2$. Hence, Theorem 1.1 can be viewed as extending the (total) dual integrality of polymatroid intersection, in the sense that it yields a $k/2$ bound on the dual integrality gap for $k$-polymatroid intersection.
Since the proof of Theorem 1.1 is constructive, it yields a $k/2$-approximation algorithm for MMSCA, which improves the current best approximation ratio $\min\{k,1+\ln n\}$.
Theorem 1.2.
There exists a $k/2$-approximation algorithm for MMSCA.
Our next result is a lower bound on the integrality gap of $\mathrm{(LP_{mono})}$, which is tight when $k$ is fixed.
Theorem 1.3.
If $\log n\geq k$, then the integrality gap of $\mathrm{(LP_{mono})}$ is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
We note that if $k$ is fixed, Theorem 1.3 implies that the integrality gap of $\mathrm{(LP_{mono})}$ is at least $k/2-\epsilon$ for any constant $\epsilon>0$, which matches the upper bound in Theorem 1.1. Another consequence is that if $k\leq\sqrt{\log n}$, then the lower bound in Theorem 1.3 satisfies $\frac{k}{2}-\frac{k(k-1)}{\log n}\geq\frac{k}{2}-1$, which almost matches the upper bound $k/2$ in Theorem 1.1.
Our proof of Theorem 1.1 is constructive as follows. We first obtain an optimal solution $(y_{1},y_{2},\ldots,y_{k})$ to $\mathrm{(LP_{mono})}$ whose supports are chains via the ellipsoid method. Then, we take the top $2/k$ of a fraction of the largest sets in the chains. We show that such a collection of sets can be partitioned into groups of $k$ sets, each of which has union $N$. The existence of this partition implies that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$. In the proof of Theorem 1.3, we construct $k$ monotone submodular functions such that the integrality gap of $\mathrm{(LP_{mono})}$ is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
Related work.
Svitkina and Tardos [18] considered the submodular facility location problem defined as follows. The input consists of a set of facilities $\mathcal{F}$, a set of demands or clients $\mathcal{D}$, a non-negative cost $c_{ij}$ to connect facility $i$ to client $j$ for each $i\in\mathcal{F}$ and $j\in\mathcal{D}$, and a facility opening cost function $f_{i}$ for each $i\in\mathcal{F}$. The goal is to assign each client to a facility so as to minimize the sum of the facility opening cost and the connection cost, where the facility opening cost for a facility $i$ is a monotone submodular function $f_{i}$ of the clients assigned to $i$. Since the sum of the facility opening cost and the connection cost for a facility $i$ is a monotone submodular function of the clients assigned to $i$, the submodular facility location problem is equivalent to MMSCA. As mentioned above, Svitkina and Tardos [18] gave a $(1+\ln|\mathcal{D}|)$-approximation algorithm for the submodular facility location problem, and showed a matching lower bound $(1-o(1))\cdot\ln|\mathcal{D}|$ when $|\mathcal{F}|$ is not fixed or bounded, while our result improves the approximation ratio to $|\mathcal{F}|/2$.
Recently, Bi et al. [3] considered the submodular multiway partition problem for monotone submodular functions, which they call monotone submodular multiway partition (Mono-Sub-MP). In Mono-Sub-MP, we are given a non-negative monotone submodular function $f:2^{N}\to\mathbb{R}_{+}$ with $f(\emptyset)=0$ and $k$ terminals $t_{1},t_{2},\ldots,t_{k}\in N$. The goal is to find a partition $S_{1},S_{2},\ldots,S_{k}$ of $N$ with $t_{i}\in S_{i}$ for every $i\in[k]$ in order to minimize $\sum_{i=1}^{k}f(S_{i})$. As the submodular multiway partition problem can be seen as a special case of MSCA, Mono-Sub-MP is a special case of MMSCA as follows. Define $N^{\prime}:=N\setminus\{t_{1},t_{2},\ldots,t_{k}\}$ and $f_{i}(S):=f(S\cup\{t_{i}\})-f(\{t_{i}\})$ for each $i\in[k]$ and $S\subseteq N^{\prime}$. Then, $f_{i}$ is a non-negative monotone submodular functions with $f_{i}(\emptyset)=0$ for every $i\in[k]$. Hence, Mono-Sub-MP can be cast as a special case of MMSCA. Bi et al. [3] showed that Mono-Sub-MP admits a $4/3$-approximation, and does not admit a $(10/9-\epsilon)$-approximation for every constant $\epsilon>0$. For further recent works related to submodular partition problems, see [1, 2, 4, 11, 14].
There are several recent works related to MSCA. Etesami [9] showed that multi-item welfare maximization under network externalities can be formulated as a special case of MSCA. Deng et al. [7] introduced the generalized load-balancing (GLB) problem, and showed that GLB is equivalent to MSCA when the outer norm is the $\ell_{1}$-norm.
Organization.
The rest of the paper is organized as follows. In Section 2, we show that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$, which yields a $k/2$-approximation algorithm for MMSCA. In Section 3, we show that the integrality gap of $\mathrm{(LP_{mono})}$ is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$ when $\log n\geq k$.
2 Approximation algorithm
In this section, we prove Theorems 1.1 and 1.2, showing that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$, which yields a $k/2$-approximation algorithm for MMSCA.
Let $N$ be a finite set of cardinality $n$, and $k\geq 2$ a positive integer. For each $i\in[k]$, let $f_{i}:2^{N}\to\mathbb{R}_{+}$ be a monotone submodular function with $f_{i}(\emptyset)=0$. Recall that $\mathrm{(LP_{mono})}$ is defined as follows:
$\textsc{$\mathrm{(LP_{mono})}$}\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
$\mathrm{subject\ to}$ $\ \ \ y_{i}\in\mathbb{R}^{2^{N}}_{+}\ \mathrm{for\ every\ }i\in[k],$
$\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)\chi_{S}=\mathbf{1}.$
To prove the upper bound of the integrality gap, we show that a $k/2$-approximate integer solution to $\mathrm{(LP_{mono})}$ can be constructed from an optimal solution to $\mathrm{(LP_{mono})}$. Since the coefficient matrix of the constraints in $\mathrm{(LP_{mono})}$ is rational, $\mathrm{(LP_{mono})}$ has a rational optimal solution $(y^{*}_{1},y^{*}_{2},\ldots,y^{*}_{k})$. Let $\mathcal{L}_{i}:=\{S\subseteq N\mid y^{*}_{i}(S)>0\}$ and $d_{i}:=|\mathcal{L}_{i}|$ for each $i\in[k]$. By the uncrossing operations, we may assume that $\mathcal{L}_{i}$ is a chain for each $i\in[k]$. Let $C_{i}^{1}\supsetneq C_{i}^{2}\supsetneq\cdots\supsetneq C_{i}^{d_{i}}$ be the members of $\mathcal{L}_{i}$ for each $i\in[k]$.
For convenience of discussion, we will convert $y^{*}_{i}$ into an integer vector $z_{i}$ by multiplying it by some constant factor, and make $z_{i}(C_{i}^{\ell})$ copies of $C_{i}^{\ell}$ for each $\ell$. To this end, we define $M$ as the product of k(k1)k(k-1) and the denominators of $y^{*}_{i}(C_{i}^{j})$ for all $i,j$. Then, $z_{i}:=M\cdot y^{*}_{i}$ is an integer vector for every $i\in[k]$. For each $i\in[k]$, we define $U_{i}^{1},U_{i}^{2},\ldots,U_{i}^{M}\subseteq N$ by setting $U_{i}^{j}:=C_{i}^{1}$ for the first $z_{i}(C_{i}^{1})$ indices $j=1,2,\ldots,z_{i}(C_{i}^{1})$, and then setting $U_{i}^{j}:=C_{i}^{2}$ for the next $z_{i}(C_{i}^{2})$ indices $j=z_{i}(C_{i}^{1})+1,z_{i}(C_{i}^{1})+2,\ldots,z_{i}(C_{i}^{1})+z_{i}(C_{i}^{2})$, and continue in this way up to $C_{i}^{d_{i}}$. For the remaining indices $j$, setting $U_{i}^{j}:=\emptyset$. In other words, for each $\ell\in[d_{i}]$, we set $U_{i}^{j}:=C_{i}^{\ell}$ for all $j$ in the following block of indices:
$1+\displaystyle\sum_{p=1}^{\ell-1}z_{i}(C_{i}^{p})\leq j\leq\displaystyle\sum_{p=1}^{\ell}z_{i}(C_{i}^{p}).$
By the definition of $U_{i}^{j}$, we have the following lemma.
Lemma 2.1.
For each $i\in[k]$ and $\ell\in[d_{i}]$, the number of $j\in[M]$ such that $U_{i}^{j}=C_{i}^{\ell}$ is $z_{i}(C_{i}^{\ell})$.
To show that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$, we will construct a partition $S_{1},S_{2},\ldots,S_{k}$ of $N$ that corresponds to a $k/2$-approximate solution to $\mathrm{(LP_{mono})}$. Such a partition can be constructed by taking the top $2/k$ fraction of the largest sets among $U_{i}^{j}$ for each $i\in[k]$, and partitioning them into groups of $k$ sets whose union is $N$. For the construction, we prepare the following two lemmas.
Lemma 2.2.
We have
$\sum_{i\in[k],j\in[M]}f_{i}(U_{i}^{j})=\sum_{S\subseteq N,i\in[k]}z_{i}(S)f_{i}(S).$
Proof.
Since $z_{i}(C_{i}^{\ell})$ is the number of $j\in[M]$ with $U_{i}^{j}=C_{i}^{\ell}$ by Lemma 2.1, we have
$\sum_{S\subseteq N,i\in[k]}z_{i}(S)f_{i}(S)=\sum_{S\in\mathcal{L}_{i},i\in[k]}z_{i}(S)f_{i}(S)=\sum_{\ell\in[d_{i}],i\in[k]}z_{i}(C_{i}^{\ell})f_{i}(C_{i}^{\ell})$
$=\sum_{\ell\in[d_{i}],i\in[k]}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\right\}\right|\cdot f_{i}(C_{i}^{\ell})$
$=\sum_{\begin{subarray}{c}i\in[k],j\in[M]\\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }\ell\in[d_{i}]\end{subarray}}f_{i}(U_{i}^{j})=\sum_{i\in[k],j\in[M]}f_{i}(U_{i}^{j}).$
The last equality follows since if $U_{i}^{j}\neq C_{i}^{\ell}$ for all $\ell=1,2,\ldots,d_{i}$, then we have $U_{i}^{j}=\emptyset$ by the definition. ∎
Lemma 2.3.
Let $a_{1},a_{2},\ldots,a_{k}\in[M]$ be positive integers such that $\sum_{i=1}^{k}a_{i}\leq M+k-1$. Then, $\bigcup_{i=1}^{k}U_{i}^{a_{i}}=N$.
Proof.
Suppose to the contrary that $\bigcup_{i=1}^{k}U_{i}^{a_{i}}\neq N$. Then, there exists $e\in N$ such that $e\notin U_{i}^{a_{i}}$ for every $i\in[k]$. Combined with $U_{i}^{1}\supseteq U_{i}^{2}\supseteq\cdots\supseteq U_{i}^{M}$ for each $i\in[k]$, this implies that $e\notin U_{i}^{j}$ for every $i\in[k]$ and integer $j$ with $a_{i}\leq j\leq M$. Hence, since $z_{i}(C_{i}^{\ell})$ is the number of $j\in[M]$ with $U_{i}^{j}=C_{i}^{\ell}$ by Lemma 2.1, we have
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}z_{i}(S)=\sum_{\begin{subarray}{c}S\in\mathcal{L}_{i},e\in S\\ i\in[k]\end{subarray}}z_{i}(S)=\sum_{\begin{subarray}{c}\ell\in[d_{i}],e\in C_{i}^{\ell}\\ i\in[k]\end{subarray}}z_{i}(C_{i}^{\ell})$
$=\sum_{\begin{subarray}{c}\ell\in[d_{i}],e\in C_{i}^{\ell}\\ i\in[k]\end{subarray}}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\right\}\right|$
$=\sum_{i\in[k]}\left|\left\{j\in[M]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }\ell\in[d_{i}]\ \mathrm{with\ }e\in C_{i}^{\ell}\right\}\right|$
$=\sum_{i\in[k]}\left|\left\{j\in[a_{i}-1]\ \middle|\ U_{i}^{j}=C_{i}^{\ell}\ \mathrm{for\ some\ }\ell\in[d_{i}]\ \mathrm{with\ }e\in C_{i}^{\ell}\right\}\right|$
$\leq\sum_{i\in[k]}(a_{i}-1)\leq M-1.$ (2.1)
Since $(y^{*}_{1},y^{*}_{2},\ldots,y^{*}_{k})$ is a feasible solution to $\mathrm{(LP_{mono})}$, we have
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}y_{i}^{*}(S)=1.$
This implies
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}z_{i}(S)=M,$
which contradicts the inequality (2). ∎
Since $M$ is divisible by k(k1)k(k-1), we can denote $M=k(k-1)\cdot m$ for some integer $m$. For $k$ positive integers $a_{1},a_{2},\ldots,a_{k}\in[M]$, we say that an ordered tuple $(a_{1},a_{2},\ldots,a_{k})$ covers $N$ if $\bigcup_{i=1}^{k}U_{i}^{a_{i}}=N$. Note that if an ordered tuple $(a_{1},a_{2},\ldots,a_{k})$ covers $N$, then there exist $S_{i}\subseteq U_{i}^{a_{i}}$ for $i=1,2,\ldots,k$ such that $(S_{1},S_{2},\ldots,S_{k})$ is a partition of $N$ and $\sum_{i=1}^{k}f_{i}(S_{i})\leq\sum_{i=1}^{k}f_{i}(U_{i}^{a_{i}})$. We first show that if there exist 2m(k1)2m(k-1) pairwise disjoint ordered tuples covering $N$, then one of them induces a partition $(S_{1},S_{2},\ldots,S_{k})$ of $N$ that corresponds to a $k/2$-approximate solution to $\mathrm{(LP_{mono})}$. We then construct 2m(k1)2m(k-1) such pairwise disjoint ordered tuples covering $N$, which completes the proof of Theorem 1.1. The following lemma is the first part of the proof:
Lemma 2.4.
Suppose that there exist 2m(k1)2m(k-1) ordered tuples
$(a_{1}^{1},\ldots,a_{k}^{1}),(a_{1}^{2},\ldots,a_{k}^{2}),\ldots,(a_{1}^{2m(k-1)},\ldots,a_{k}^{2m(k-1)})$
such that
  • (1)
    $a_{i}^{j}\in[M]$ for every $i\in[k]$ and $j\in[2m(k-1)]$,
  • (2)
    $a_{i}^{j}\neq a_{i}^{\ell}$ for every $i\in[k]$ and $j,\ell\in[2m(k-1)]$ with $j\neq\ell$,
  • (3)
    every ordered tuple covers $N$.
Then, the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$.
Proof.
By Lemma 2.2, we have
$\sum_{S\subseteq N,i\in[k]}z_{i}(S)f_{i}(S)=\sum_{i\in[k],j\in[M]}f_{i}(U_{i}^{j})\geq\sum_{i\in[k],j\in[2m(k-1)]}f_{i}(U_{i}^{a_{i}^{j}})$
$\geq 2m(k-1)\cdot\min\left\{\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}})\ \middle|\ j\in[2m(k-1)]\right\}.$ (2.2)
The first inequality follows since by the monotonicity of $f_{i}$, we have $f_{i}(U_{i}^{j})\geq f_{i}(\emptyset)=0$ for every $i\in[k]$ and $j\in[M]$. Let
$j^{*}:=\underset{j\in[2m(k-1)]}{\operatorname{argmin}}\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}}).$
Then, by (2) we have
$\sum_{S\subseteq N,i\in[k]}z_{i}(S)f_{i}(S)\geq 2m(k-1)\cdot\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j^{*}}}).$
Since $z_{i}=M\cdot y^{*}_{i}=mk(k-1)\cdot y^{*}_{i}$ for each $i\in[k]$, this implies
$\dfrac{k}{2}\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\geq\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j^{*}}}).$ (2.3)
Since $\bigcup_{i=1}^{k}U_{i}^{a_{i}^{j^{*}}}=N$, there exists a partition $S_{1},S_{2},\ldots,S_{k}$ of $N$ such that $S_{i}\subseteq U_{i}^{a_{i}^{j^{*}}}$ for every $i\in[k]$. Then, by (2.3) and the monotonicity of $f_{i}$, we have
$\dfrac{k}{2}\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\geq\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j^{*}}})\geq\sum_{i\in[k]}f_{i}(S_{i}).$
This implies that the integrality gap of $\mathrm{(LP_{mono})}$ is at most $k/2$. ∎
We now prove Theorem 1.1 by constructing 2m(k1)2m(k-1) pairwise disjoint ordered tuples covering $N$, using only the smallest $2/k$ fraction of the indices $j=1,2,\ldots,M$ of $U_{i}^{j}$ for each $i\in[k]$.
Proof of Theorem 1.1.
By Lemma 2.4, it suffices to show that there exist 2m(k1)2m(k-1) ordered tuples satisfying the conditions in Lemma 2.4. For each positive integer $a\in\mathbb{Z}_{>0}$, define $\mathrm{mod}(a)$ to be the unique integer $r\in[2m(k-1)]$ such that $a\equiv r\pmod{2m(k-1)}$. For each $i\in[k]$ and $j\in[2m(k-1)]$, define $a_{i}^{j}$ as follows:
$a_{i}^{j}:=\begin{cases}\mathrm{mod}(2m(i-1)+j)&\mathrm{if\ }1\leq i\leq k-1,\\ \mathrm{mod}((2m-j+1)(k-1))-\left\lfloor\dfrac{j-1}{2m}\right\rfloor&\mathrm{if\ }i=k.\end{cases}$
We show that $(a_{i}^{j})_{i\in[k],j\in[2m(k-1)]}$ satisfies conditions (1)–(3) in Lemma 2.4. For condition (1), we show that $a_{i}^{j}\in[M]$ for each $i\in[k]$ and $j\in[2m(k-1)]$. If $1\leq i\leq k-1$, then we have
$1\leq a_{i}^{j}=\mathrm{mod}(2m(i-1)+j)\leq 2m(k-1)\leq km(k-1)=M.$
If $i=k$, then we have
$a_{i}^{j}=\mathrm{mod}((2m-j+1)(k-1))-\left\lfloor\dfrac{j-1}{2m}\right\rfloor.$
Since (2mj+1)(k1)(2m-j+1)(k-1) is a multiple of k1k-1, we have
$\mathrm{mod}((2m-j+1)(k-1))=p(k-1)$
for some positive integer $p$. This implies
$\mathrm{mod}((2m-j+1)(k-1))\geq k-1.$
Hence, we have
$1=k-1-(k-2)=k-1-\left\lfloor\dfrac{2m(k-1)-1}{2m}\right\rfloor\leq\mathrm{mod}((2m-j+1)(k-1))-\left\lfloor\dfrac{j-1}{2m}\right\rfloor,$
and
$\mathrm{mod}((2m-j+1)(k-1))-\left\lfloor\dfrac{j-1}{2m}\right\rfloor\leq 2m(k-1)\leq km(k-1)=M.$
Therefore, condition (1) holds for $(a_{i}^{j})_{i\in[k],j\in[2m(k-1)]}$.
For condition (2), we next show that $a_{i}^{j}\neq a_{i}^{\ell}$ for each $i\in[k]$ and $j,\ell\in[2m(k-1)]$ with $j\neq\ell$. If $1\leq i\leq k-1$, then by the definition of $a_{i}^{j}$, it is clear that $a_{i}^{j}\neq a_{i}^{\ell}$ holds for each $j,\ell\in[2m(k-1)]$ with $j\neq\ell$. Consider the case when $i=k$. Suppose to the contrary that $a_{k}^{j}=a_{k}^{\ell}$ holds for some $j,\ell\in[2m(k-1)]$ with $j\neq\ell$. This implies that
$\mathrm{mod}((2m-j+1)(k-1))-\left\lfloor\dfrac{j-1}{2m}\right\rfloor=\mathrm{mod}((2m-\ell+1)(k-1))-\left\lfloor\dfrac{\ell-1}{2m}\right\rfloor.$ (2.4)
Since $\mathrm{mod}((2m-j+1)(k-1))$ and $\mathrm{mod}((2m-\ell+1)(k-1))$ are multiples of k1k-1, it follows that
$\left\lfloor\dfrac{j-1}{2m}\right\rfloor-\left\lfloor\dfrac{\ell-1}{2m}\right\rfloor$
is a multiple of k1k-1 by (2.4). Combined with
$0\leq\left\lfloor\dfrac{j-1}{2m}\right\rfloor,\left\lfloor\dfrac{\ell-1}{2m}\right\rfloor\leq k-2,$
this implies that
$\left\lfloor\dfrac{j-1}{2m}\right\rfloor=\left\lfloor\dfrac{\ell-1}{2m}\right\rfloor.$
Hence, we have $2m(p-1)+1\leq j,\ell\leq 2mp$ for some $p\in[k-1]$. Let $j^{\prime}:=2mp-j$ and $\ell^{\prime}:=2mp$-$\ell$. By the definition of $p$, we have $0\leq j^{\prime},\ell^{\prime}\leq 2m-1$. By (2.4), we have
$\mathrm{mod}((2m-j+1)(k-1))=\mathrm{mod}((2m-\ell+1)(k-1)).$ (2.5)
We also have
$\mathrm{mod}((2m-j+1)(k-1))=\mathrm{mod}((2m-2mp+j^{\prime}+1)(k-1))=(j^{\prime}+1)(k-1),$
$\mathrm{mod}((2m-\ell+1)(k-1))=\mathrm{mod}((2m-2mp+\ell^{\prime}+1)(k-1))=(\ell^{\prime}+1)(k-1).$ (2.6)
Hence, by (2.5) and (2) we have
$(j^{\prime}+1)(k-1)=(\ell^{\prime}+1)(k-1).$
This implies $j^{\prime}=\ell^{\prime}$, which contradicts $j\neq\ell$.
For condition (3), we show that $\bigcup_{i=1}^{k}U_{i}^{a_{i}^{j}}=N$ holds for every $j\in[2m(k-1)]$. By Lemma 2.3, it suffices to show that $\sum_{i=1}^{k}a_{i}^{j}\leq M+k-1$ for every $j\in[2m(k-1)]$. Take any $j\in[2m(k-1)]$. Let $p\in[k-1]$ be the unique positive integer such that $2m(p-1)+1\leq j\leq 2mp$. Let $j^{\prime}:=2mp-j$. Then, we have $0\leq j^{\prime}\leq 2m-1$. By the definition of $a_{i}^{j}$, we have
$\sum_{i=1}^{k}a_{i}^{j}=\sum_{i=1}^{k-1}\mathrm{mod}(2m(i-1)+j)+\mathrm{mod}((2m-j+1)(k-1))-\left\lfloor\dfrac{j-1}{2m}\right\rfloor$
$=\sum_{i=1}^{k-1}\mathrm{mod}(2m(i-1)+2mp-j^{\prime})+\mathrm{mod}((2m-2mp+j^{\prime}+1)(k-1))-\left\lfloor\dfrac{2mp-j^{\prime}-1}{2m}\right\rfloor$
$=\sum_{i=1}^{k-p}(2m(i-1)+2mp-j^{\prime})+\sum_{i=k-p+1}^{k-1}(2m(i-1)+2mp-j^{\prime}-2m(k-1))$
$+(j^{\prime}+1)(k-1)-(p-1)$
$=\sum_{i=1}^{k-1}(2m(i-1)+2mp-j^{\prime})-2m(k-1)(p-1)+(j^{\prime}+1)(k-1)-(p-1)$
$=mk(k-1)-(k-1)(2m-2mp+j^{\prime})-2m(k-1)(p-1)+(j^{\prime}+1)(k-1)-(p-1)$
$=M+k-p\leq M+k-1.$
This completes the proof. ∎
The proofs of Lemma 2.4 and Theorem 1.1 provide the construction of a partition $S_{1},S_{2},\ldots,S_{k}$ of $N$ that corresponds to a $k/2$-approximate solution to $\mathrm{(LP_{mono})}$, and hence it is also a $k/2$-approximate solution to MMSCA. The construction works as follows. We first compute a rational optimal solution $(y^{*}_{1},y^{*}_{2},\ldots,y^{*}_{k})$ to $\mathrm{(LP_{mono})}$ whose supports are chains, and define $U_{i}^{j}$ for each $i\in[k]$ and $j\in[M]$. We then define 2m(k1)2m(k-1) ordered tuples $(a_{1}^{j},a_{2}^{j},\ldots,a_{k}^{j})$ covering $N$ as in the proof of Theorem 1.1, and compute the index $j^{*}$ minimizing $\sum_{i=1}^{k}f_{i}(U_{i}^{a_{i}^{j}})$ in the proof of Lemma 2.4. Finally, we obtain sets $U_{1}^{a_{1}^{j^{*}}},U_{2}^{a_{2}^{j^{*}}},\ldots,U_{k}^{a_{k}^{j^{*}}}$, which induces a desired partition $S_{1},S_{2},\ldots,S_{k}$ of $N$.
We prove Theorem 1.2 by showing that the above construction can be done in polynomial time.
Proof of Theorem 1.2.
We can obtain a rational optimal solution $(y_{1}^{*},y_{2}^{*},\ldots,y_{k}^{*})$ to $\mathrm{(LP_{mono})}$ whose supports $\mathcal{L}_{1},\mathcal{L}_{2},\ldots,\mathcal{L}_{k}$ are chains in polynomial time via the ellipsoid method [5] (see Appendix A for details). Then, we define $U_{i}^{j}$ for each $i\in[k]$ and $j\in[M]$, and $a_{i}^{j}$ for each $i\in[k]$ and $j\in[2m(k-1)]$ as in the proof of Theorem 1.1. If we can compute
$j^{*}\in\underset{j\in[2m(k-1)]}{\operatorname{argmin}}\sum_{i=1}^{k}f_{i}(U_{i}^{a_{i}^{j}}),$
then any partition $S_{1},S_{2},\ldots,S_{k}$ of $N$ such that $S_{i}\subseteq U_{i}^{a_{i}^{j^{*}}}$ for every $i\in[k]$ satisfies
$\sum_{i=1}^{k}f_{i}(S_{i})\leq\sum_{i=1}^{k}f_{i}(U_{i}^{a_{i}^{j^{*}}})\leq\dfrac{k}{2}\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)$
$\leq\dfrac{k}{2}\cdot\min\left\{\sum_{i=1}^{k}f_{i}(T_{i})\ \middle|\ T_{1},T_{2},\ldots,T_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\},$
as desired. Hence, it suffices to show that $j^{*}$ can be computed in polynomial time. To find $j^{*}$, we set $j_{1}:=1$. Then, for q=2,3,,q=2,3,..., we define $j_{q}$ as the minimum integer such that $j_{q}>j_{q-1}$ and there exists $i\in[k]$ with $U_{i}^{a_{i}^{j_{q}}}\neq U_{i}^{a_{i}^{j_{q-1}}}$. Recall that $n=|N|$. Since for each $i\in[k]$, we have $U_{i}^{1}\supseteq U_{i}^{2}\supseteq\cdots\supseteq U_{i}^{M}$ and there are at most $|\mathcal{L}_{i}|\leq n$ distinct sets among $\{U_{i}^{j}\}_{j=1}^{M}$, we only need to define $j_{q}$ for $q=1,2,\ldots,Q$ with $Q\leq kn$. Since we have
$j^{*}\in\underset{j\in\{j_{1},j_{2},\ldots,j_{Q}\}}{\operatorname{argmin}}\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j}}),$
we can find $j^{*}$ by computing the values of $\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j_{q}}})$ for all $q\in[Q]$. By the definition of $U_{i}^{j}$ and $a_{i}^{j}$, we can compute $j_{q}$ in polynomial time for each $q\in[Q]$. Hence, since $Q\leq kn$, we can compute $j_{q}$ and $\sum_{i\in[k]}f_{i}(U_{i}^{a_{i}^{j_{q}}})$ for all $q\in[Q]$ in polynomial time. ∎
3 Lower bound on the integrality gap
In this section, we show that the integrality gap of $\mathrm{(LP_{mono})}$ is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$ when $\log n\geq k$. To this end, we explicitly construct monotone submodular functions $f_{1},f_{2},\ldots,f_{k}$ such that the integrality gap of $\mathrm{(LP_{mono})}$ is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$.
Suppose that $\log n\geq k$. Let $p$ be the unique positive integer such that $pk\leq\lfloor\log n\rfloor\leq(p+1)k-1$. We define
$\mathcal{V}:=\left\{v\in\mathbb{Z}_{+}^{k}\ \middle|\ \sum_{i=1}^{k}v_{i}=pk-k+1\right\}.$
By stars and bars (see e.g., [17, Chapter 1]), we have
$|\mathcal{V}|=\binom{pk}{k-1}\leq 2^{pk}\leq 2^{\lfloor\log n\rfloor}\leq n=|N|.$
Choose $e_{v}\in N$ for each $v\in\mathcal{V}$ so that all $e_{v}$ are distinct. For each $i\in[k]$, let
$N_{i}:=\{e_{v}\in N\mid v\in\mathcal{V},v_{i}\geq 1\},$
$Z_{i}:=\{e_{v}\in N\mid v\in\mathcal{V},v_{i}=0\}.$
For each $i\in[k]$, define $f_{i}:2^{N}\to\mathbb{R}$ as follows: for all $S\subseteq N$,
$f_{i}(S):=\begin{cases}\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in S\cap N_{i}\}\}&\mathrm{if\ }S\cap N_{i}\neq\emptyset\ \mathrm{and\ }S\cap Z_{i}=\emptyset,\\ 0&\mathrm{if\ }S\cap N_{i}=\emptyset\ \mathrm{and\ }S\cap Z_{i}=\emptyset,\\ 2pk+1&\mathrm{if\ }S\cap Z_{i}\neq\emptyset.\end{cases}$
Lemma 3.1.
$f_{i}$ is monotone submodular for each $i\in[k]$.
Proof.
We first show monotonicity of $f_{i}$. Take any $S,T\subseteq N$ with $S\subseteq T$. If $S\cap N_{i}=S\cap Z_{i}=\emptyset$, then we have $f_{i}(S)=0\leq f_{i}(T)$. If $S\cap Z_{i}\neq\emptyset$, then we have $T\cap Z_{i}\neq\emptyset$, which implies $f_{i}(S)=f_{i}(T)=2pk+1$. Consider the case when $S\cap N_{i}\neq\emptyset$ and $S\cap Z_{i}=\emptyset$. If $T\cap Z_{i}\neq\emptyset$, then we have
$f_{i}(S)=\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in S\cap N_{i}\}\}\leq 2p+1-1=2p\leq 2pk+1=f_{i}(T).$
If $T\cap Z_{i}=\emptyset$, then since $T\cap N_{i}\neq\emptyset$ follows from $S\cap N_{i}\neq\emptyset$, we have
$f_{i}(S)$ $=\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in S\cap N_{i}\}\}$
$\leq\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in T\cap N_{i}\}\}=f_{i}(T).$
Therefore, $f_{i}$ is monotone.
We next show that $f_{i}$ is submodular. Take any $S,T\subseteq N$. If $S\cap Z_{i}\neq\emptyset$, then by the monotonicity of $f_{i}$, we have
$f_{i}(S)+f_{i}(T)=2pk+1+f_{i}(T)\geq 2pk+1+f_{i}(S\cap T)\geq f_{i}(S\cup T)+f_{i}(S\cap T).$
The last inequality follows since $f_{i}(X)\leq 2pk+1$ holds for every $X\subseteq N$ by the definition of $f_{i}$. If $S\cap N_{i}=S\cap Z_{i}=\emptyset$, then since $(S\cup T)\cap N_{i}=T\cap N_{i}$ and $(S\cup T)\cap Z_{i}=T\cap Z_{i}$, we have $f_{i}(S\cup T)=f_{i}(T)$ by the definition of $f_{i}$. This implies that
$f_{i}(S)+f_{i}(T)=f_{i}(S)+f_{i}(S\cup T)\geq f_{i}(S\cup T)+f_{i}(S\cap T).$
The last inequality follows from the monotonicity of $f_{i}$. Hence, by the symmetry between $S$ and $T$, the remaining case is when $S\cap N_{i}\neq\emptyset$ and $S\cap Z_{i}=\emptyset$, and $T\cap N_{i}\neq\emptyset$ and $T\cap Z_{i}=\emptyset$. In this case, we have $(S\cup T)\cap N_{i}\neq\emptyset$ and $(S\cup T)\cap Z_{i}=\emptyset$, which implies that
$f_{i}(S\cup T)=\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in(S\cup T)\cap N_{i}\}\}.$
Take $e_{v^{*}}\in\mathrm{argmax}\{2p+1-v_{i}\mid e_{v}\in(S\cup T)\cap N_{i}\}$ with the corresponding vector $v^{*}\in\mathcal{V}$. Then, we have $e_{v^{*}}\in S\cap N_{i}$ or $e_{v^{*}}\in T\cap N_{i}$. We may assume without loss of generality that $e_{v^{*}}\in S\cap N_{i}$. Then, we have $e_{v^{*}}\in\mathrm{argmax}\{2p+1-v_{i}\mid e_{v}\in S\cap N_{i}\}$. Hence, we have
$f_{i}(S)+f_{i}(T)=\max\{0,2p+1-v^{*}_{i}\}+f_{i}(T)=f_{i}(S\cup T)+f_{i}(T)$ $\geq f_{i}(S\cup T)+f_{i}(S\cap T).$
The last inequality follows from the monotonicity of $f_{i}$. Therefore, $f_{i}$ is submodular. ∎
We now show that for the monotone submodular functions $f_{1},f_{2},\ldots,f_{k}$ defined above, any optimal solution $(y_{1}^{*},\ldots,y_{k}^{*})$ to $\mathrm{(LP_{mono})}$ satisfies
$\left(\frac{k}{2}-\frac{k(k-1)}{\log n}\right)\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)$
$\leq\min\left\{\sum_{i=1}^{k}f_{i}(S_{i})\ \middle|\ S_{1},S_{2},\ldots,S_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$ (3.1)
The inequality (3) implies that the integrality gap of $\mathrm{(LP_{mono})}$ is at least $\frac{k}{2}-\frac{k(k-1)}{\log n}$, which completes the proof of Theorem 1.3.
To prove (3), we first give an upper bound on the value $\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)$. To this end, we construct a feasible solution to $\mathrm{(LP_{mono})}$. For each $i\in[k]$ and $j\in[pk-k+1]$, let $N_{i}^{j}:=\{e_{v}\in N\mid v\in\mathcal{V},v_{i}\geq j\}$. Note that $N_{i}^{1}=N_{i}$. In what follows, we will use $N_{i}^{1}$ instead of $N_{i}$ for convenience. Let $N^{c}:=N\setminus\{e_{v}\in N\mid v\in\mathcal{V}\}$. For each $i\in[k]$, define $y^{\prime}_{i}\in\mathbb{R}_{+}^{2^{N}}$ as follows: for all $S\subseteq N$,
$y^{\prime}_{i}(S):=\begin{cases}\dfrac{1}{pk-k+1}&\mathrm{if\ }i=1\ \mathrm{and\ }\vskip 3.0ptS=N_{1}^{j}\cup N^{c}\ \mathrm{for\ some\ }j\in[pk-k+1],\\ \vskip 3.0pt\dfrac{1}{pk-k+1}&\mathrm{if\ }2\leq i\leq k\ \mathrm{and\ }S=N_{i}^{j}\ \mathrm{for\ some\ }j\in[pk-k+1],\\ 0&\mathrm{otherwise}.\end{cases}$
Lemma 3.2.
$(y^{\prime}_{1},y^{\prime}_{2},\ldots,y^{\prime}_{k})$ is a feasible solution to $\mathrm{(LP_{mono})}$.
Proof.
To show the feasibility of $(y^{\prime}_{1},y^{\prime}_{2},\ldots,y^{\prime}_{k})$, we show that
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}y^{\prime}_{i}(S)=1$ (3.2)
for every $e\in N$. For any $e\in N^{c}$, we have
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}y^{\prime}_{i}(S)=\sum_{j\in[pk-k+1]}y^{\prime}_{1}(N_{1}^{j}\cup N^{c})=\sum_{j\in[pk-k+1]}\dfrac{1}{pk-k+1}=1.$
Take any $e\in N\setminus N^{c}$. Since $e=e_{v}$ for some $v\in\mathcal{V}$, we have
$\sum_{\begin{subarray}{c}S\subseteq N,e\in S\\ i\in[k]\end{subarray}}y^{\prime}_{i}(S)$ $=\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ v_{1}\geq j\end{subarray}}y^{\prime}_{1}(N_{1}^{j}\cup N^{c})+\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ 2\leq i\leq k,v_{i}\geq j\end{subarray}}y^{\prime}_{i}(N_{i}^{j})$
$=\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ v_{1}\geq j\end{subarray}}\frac{1}{pk-k+1}+\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ 2\leq i\leq k,v_{i}\geq j\end{subarray}}\frac{1}{pk-k+1}$
$=\sum_{\begin{subarray}{c}j\in[pk-k+1]\\ i\in[k],v_{i}\geq j\end{subarray}}\frac{1}{pk-k+1}$
$=\frac{pk-k+1}{pk-k+1}=1.$
The second last equality follows since $\sum_{i=1}^{k}v_{i}=pk-k+1$. Hence, (3.2) holds for every $e\in N$, which implies that $(y^{\prime}_{1},y^{\prime}_{2},\ldots,y^{\prime}_{k})$ is a feasible solution to $\mathrm{(LP_{mono})}$. ∎
By Lemma 3.2 and the optimality of $y^{*}_{1},y^{*}_{2},\ldots,y^{*}_{k}$, we have
$\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\leq\sum_{S\subseteq N,i\in[k]}y^{\prime}_{i}(S)f_{i}(S)$
$=\sum_{j\in[pk-k+1]}\frac{1}{pk-k+1}f_{1}(N_{1}^{j}\cup N^{c})+\sum_{2\leq i\leq k,j\in[pk-k+1]}\frac{1}{pk-k+1}f_{i}(N_{i}^{j})$
$=\sum_{j\in[pk-k+1]}\frac{1}{pk-k+1}\cdot\max\{0,2p+1-j\}+\sum_{2\leq i\leq k,j\in[pk-k+1]}\frac{1}{pk-k+1}\cdot\max\{0,2p+1-j\}$
$=\frac{pk(2p+1)}{pk-k+1}.$ (3.3)
To prove (3), we next provide a lower bound on the right hand side of (3).
Lemma 3.3.
Let $S_{1},S_{2},\ldots,S_{k}$ be any partition of $N$. Then, we have
$\sum_{i=1}^{k}f_{i}(S_{i})\geq pk+k.$
Proof.
If $S_{j}\cap Z_{j}\neq\emptyset$ for some $j\in[k]$, then we have
$\sum_{i=1}^{k}f_{i}(S_{i})\geq f_{j}(S_{j})=2pk+1\geq pk+k.$
Hence, it suffices to consider the case when $S_{i}\cap Z_{i}=\emptyset$ for every $i\in[k]$. For each $i\in[k]$, define $x^{i}\in\mathcal{V}$ as follows:
$x^{i}_{j}:=\begin{cases}pk-k+1&\mathrm{if\ }j=i,\\ 0&\mathrm{if\ }j\in[k]\mathrm{\ and\ }j\neq i.\end{cases}$
Then, since $S_{i}\cap Z_{i}=\emptyset$ for each $i\in[k]$, we have $e_{x^{i}}\in S_{i}$ for every $i\in[k]$. Hence, we have $S_{i}\cap N_{i}\neq\emptyset$ for every $i\in[k]$. For each $i\in[k]$, we define $b_{i}:=\min\{v_{i}\mid e_{v}\in S_{i}\cap N_{i}\}$. We now show that $\sum_{i=1}^{k}b_{i}\leq pk$. Suppose to the contrary that $\sum_{i=1}^{k}b_{i}\geq pk+1$. Then, since $\sum_{i=1}^{k}(b_{i}-1)\geq pk-k+1$, there exists $v\in\mathcal{V}$ such that $v_{i}\leq b_{i}-1$ for every $i\in[k]$. Choose $i\in[k]$ such that $e_{v}\in S_{i}$. Since $S_{i}\cap Z_{i}=\emptyset$ by the assumption, we have $e_{v}\in S_{i}\cap N_{i}$, which contradicts the definition of $b_{i}$. Hence, we have $\sum_{i=1}^{k}b_{i}\leq pk$. Therefore, since $S_{i}\cap N_{i}\neq\emptyset$ and $S_{i}\cap Z_{i}=\emptyset$ for each $i\in[k]$, we have
$\sum_{i=1}^{k}f_{i}(S_{i})=\sum_{i=1}^{k}\max\{0,\max\{2p+1-v_{i}\mid e_{v}\in S_{i}\cap N_{i}\}\}$
$\geq\sum_{i=1}^{k}\max\{2p+1-v_{i}\mid e_{v}\in S_{i}\cap N_{i}\}=2pk+k-\sum_{i=1}^{k}\min\{v_{i}\mid e_{v}\in S_{i}\cap N_{i}\}$
$=2pk+k-\sum_{i=1}^{k}b_{i}\geq 2pk+k-pk=pk+k.$
By Lemma 3.3, we have
$\min\left\{\sum_{i=1}^{k}f_{i}(S_{i})\ \middle|\ S_{1},S_{2},\ldots,S_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}$
$\geq pk+k=\frac{2k(p+1)^{2}(pk-k+1)}{2(p+1)(pk-k+1)}$
$=\frac{2k(p+1)(p^{2}k-k+p+1)}{2(p+1)(pk-k+1)}=\frac{2k(p+1)(p^{2}k+(p-1)(k-1)-pk+2p)}{2(p+1)(pk-k+1)}$
$\geq\frac{2k(p+1)(p^{2}k-pk+2p)}{2(p+1)(pk-k+1)}=\frac{pk(2p+2)(pk-k+2)}{2(p+1)(pk-k+1)}$
$\geq\frac{pk(2p+1)(pk-k+2)}{2(p+1)(pk-k+1)}=\frac{pk(2p+1)}{pk-k+1}\cdot\left(\frac{k}{2}-\frac{k(k-1)}{k(p+1)}\right)$
$\geq\frac{pk(2p+1)}{pk-k+1}\cdot\left(\frac{k}{2}-\frac{k(k-1)}{\log n}\right).$ (3.4)
The last inequality follows from the definition of $p$. Combining (3) and (3), we have
$\left(\frac{k}{2}-\frac{k(k-1)}{\log n}\right)\cdot\sum_{S\subseteq N,i\in[k]}y^{*}_{i}(S)f_{i}(S)\leq\left(\frac{k}{2}-\frac{k(k-1)}{\log n}\right)\cdot\frac{pk(2p+1)}{pk-k+1}$
$\leq\min\left\{\sum_{i=1}^{k}f_{i}(S_{i})\ \middle|\ S_{1},S_{2},\ldots,S_{k}\ \mathrm{is\ a\ partition\ of\ }N\right\}.$
Thus, the inequality (3) holds, which completes the proof of Theorem 1.3.
Acknowledgements
This work was supported by JST ERATO Grant Number JPMJER2301, Japan.
Appendix A Convex program based on the Lovász extension
In this section, we show that (LP) is equivalent to the convex program relaxation of MSCA introduced by Chekuri and Ene [5]. Let $N:=\{1,\ldots,n\}$. Let $f_{i}:2^{N}\to\mathbb{R}$ be a submodular function with $f_{i}(\emptyset)=0$ for each $i\in[k]$. Recall that (LP) is defined as follows:
$\textsc{(LP)}\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)$
$\mathrm{subject\ to}$ $\ \ \ y_{i}\in\mathbb{R}^{2^{N}}_{+}\ \mathrm{for\ every\ }i\in[k],$
$\ \ \ \sum_{S\subseteq N,i\in[k]}y_{i}(S)\chi_{S}=\mathbf{1}.$
For a set function $f:2^{N}\to\mathbb{R}$ with $f(\emptyset)=0$, the Lovász extension $\hat{f}:[0,1]^{n}\to\mathbb{R}$ of $f$ is defined as follows: Take any $x\in[0,1]^{n}$. Relabel the elements in $N$ so that $x_{1}\geq x_{2}\geq\cdots\geq x_{n}$. For convenience, let $x_{n+1}:=0$. Let $S_{i}:=\{1,2,\ldots,i\}$ for each $i\in[n]$. The value of $\hat{f}(x)$ is defined as follows:
$\hat{f}(x):=\sum_{i=1}^{n}(x_{i}-x_{i+1})f(S_{i}).$
Chekuri and Ene [5] introduced the following convex program relaxation of MSCA based on the Lovász extension:
$(\textsc{LE-Rel})\ \ \ \ \ \mathrm{minimize}$ $\ \ \ \sum_{i=1}^{k}\hat{f}_{i}(x^{i})$
$\mathrm{subject\ to}$ $\ \ \ x^{i}\in[0,1]^{n}\ \mathrm{for\ every\ }i\in[k],$
$\ \ \ \sum_{i=1}^{k}x^{i}_{j}=1\ \mathrm{for\ every\ }j\in[n].$
We now show that any feasible solution $({x^{1}},\ldots,x^{k})$ to (LE-Rel) can be translated into a feasible solution $(y_{1},y_{2},\ldots,y_{k})$ to (LP) that has the same objective value and whose supports are chains. Take any feasible solution $({x^{1}},\ldots,x^{k})$ to (LE-Rel). We will describe a translation from $x^{i}$ to $y_{i}$ for each $i\in[k]$. We first relabel the elements in $N$ so that $x^{i}_{1}\geq x^{i}_{2}\geq\cdots\geq x^{i}_{n}$. Let $x^{i}_{n+1}:=0$ and $S_{j}:=\{1,2,\ldots,j\}$ for each $j\in[n]$. Then, for all $S\subseteq N$ define $y_{i}(S)$ as follows:
$y_{i}(S):=\begin{cases}x_{j}^{i}-x_{j+1}^{i}&\mathrm{if\ }S=S_{j}\ \mathrm{and\ }x_{j}^{i}-x_{j+1}^{i}>0\ \mathrm{for\ some\ }j\in[n],\\ 0&\mathrm{otherwise}.\end{cases}$
Then, $(y_{1},y_{2},\ldots,y_{k})$ is a feasible solution to (LP) because for each $i\in[k]$ we have
$\sum_{S\subseteq N}y_{i}(S)\chi_{S}=\sum_{j\in[n]}(x^{i}_{j}-x^{i}_{j+1})\chi_{S_{j}}=x^{i}$
and
$\sum_{S\subseteq N,i\in[k]}y_{i}(S)\chi_{S}=\sum_{i\in[k]}x^{i}=\mathbf{1}.$
For each $i\in[k]$, we also have
$\sum_{S\subseteq N}y_{i}(S)f_{i}(S)=\sum_{j\in[n]}(x_{j}^{i}-x_{j+1}^{i})f_{i}(S_{j})=\hat{f_{i}}(x^{i}).$
Hence, we obtain
$\sum_{S\subseteq N,i\in[k]}y_{i}(S)f_{i}(S)=\sum_{i=1}^{k}\hat{f_{i}}(x^{i}),$
which implies that $(y_{1},y_{2},\ldots,y_{k})$ has the same objective value as $(x^{1},\ldots,x^{k})$. Similarly, we can translate any feasible solution to (LP) whose supports are chains into a feasible solution to (LE-Rel) that has the same objective value.
Since (LE-Rel) can be solved in polynomial time via the ellipsoid method [5], an optimal solution to (LP) whose supports are chains can be obtained in polynomial time. Chekuri and Ene [6] provided another equivalent linear programming formulation of MSCA called LE-Rel-Primal. We note that (LP) is equivalent to LE-Rel-Primal.
Appendix B $k$-polymatroid intersection
Let $N$ be a finite set and $k\geq 2$ a positive integer. For each $i\in[k]$, let $f_{i}:2^{N}\to\mathbb{R}$ be a submodular function with $f_{i}(\emptyset)=0$. For a set function $f:2^{N}\to\mathbb{R}$, we define $P(f):=\{x\in\mathbb{R}^{N}\mid x(S)\leq f(S)\ \mathrm{for\ every\ }S\subseteq N\}$, where $x(S):=\sum_{e\in S}x(e)$ for every $S\subseteq N$. Let $w\in\mathbb{R}_{+}^{N}$ be a weight vector. The weighted $k$-polymatroid intersection problem is defined as follows:
$(\mathrm{LP}_{k\text{-poly}})\ \ \ \ \ \mathrm{maximize}$ $\ \ \ w^{\top}x$
$\mathrm{subject\ to}$ $\ \ \ x\in P(f_{i})\ \mathrm{for\ every\ }i\in[k].$
The linear program $\mathrm{LP}_{k\text{-poly}}$ corresponds to the weighted polymatroid intersection when $k=2$. When $w=\mathbf{1}$, the linear program $\mathrm{LP}_{k\text{-poly}}$ is the dual linear program of (LP). For the case when $f_{1},f_{2},\ldots,f_{k}$ are matroid rank functions (i.e., the case of weighted $k$-matroid intersection), the integrality gap of $\mathrm{LP}_{k\text{-poly}}$ is $2$ when $k=3$ [13], and it equals k1k-1 for arbitrary $k$ in the unweighted case $w=\mathbf{1}$ [12].